### v1

# TaCo: Targeted Concept Removal in Output Embeddings for NLP via Information Theory and Explainability

Abstract
The fairness of Natural Language Processing (NLP) models has emerged as a crucial concern. Information theory indicates that to achieve fairness, a model should not be able to predict sensitive variables, such as gender, ethnicity, and age. However, information related to these variables often appears implicitly in language, posing a challenge in identifying and mitigating biases effectively.
To tackle this issue, we present a novel approach that operates at the embedding level of an NLP model, independent of the specific architecture. Our method leverages insights from recent advances in XAI techniques and employs an embedding transformation to eliminate implicit information from a selected variable. By directly manipulating the embeddings in the final layer, our approach enables a seamless integration into existing models without requiring significant modifications or retraining.
In evaluation, we show that the proposed post-hoc approach significantly reduces gender-related associations in NLP models while preserving the overall performance and functionality of the models.
An implementation of our method is available: https://github.com/fanny-jourdan/TaCo
1 Introduction
The multiple and impressive successes of Artificial Intelligence and Machine Learning have made them more and more ubiquitous in applications where a model’s decisions has genuine repercussions on people’s lives. These decisions, however, typically come with biases that can lead to unfair, unethical consequences. To guard against such consequences, we must identify and understand the biased information these models utilize. This links the task of ensuring an ethical comportment of a model with its interpretability or explainability: if we can show the real reasons or features why a model makes the predictions or recommendations it does, then we can control for biases among these features.
In this paper, we propose a method for finding factors that influence model predictions using an algebraic decomposition of the latent representations of the model into orthogonal dimensions. Using the method of Singular Value Decomposition (SVD), which is well suited to our particular test case, we isolate various dimensions that contribute to a prediction. As our case study, we study the use of gender information in predicting occupations in the Bios dataset De-Arteaga et al. (2019) by a RoBERTa model Liu et al. (2019a). By intervening on the latent representations to remove certain dimensions and running the model on the edited representations, we show first a causal connection between various dimensions and model predictions about gender and separately about occupation. This enables us then to eliminate dimensions in the decomposition that convey gender bias in the decision without much affecting the model’s predictive performance on occupations. Because SVD gives us orthogonal dimensions, we know that once we have removed a dimension, the information it contains is no longer in the model.
Crucially, we show how to link various dimensions in the SVD decomposition of the latent representation with concepts like gender and its linguistic manifestations that matter to fairness, providing not only a faithful explanation of the model’s behaviour but a humanly understandable or plausible one as well. Faithfulness provides a causal connection and produces factors necessary and sufficient for the model to produce its prediction given a certain input. Plausibility has to do with how acceptable the purported explanation is to humans Jacovi and Goldberg (2020).
In addition, we demonstrate that our method can reach close to the theoretically ideal point of minimizing algorithmic bias while preserving strong predictive accuracy, and we offer a way of quantifying various operations on representations. Our method also has a relatively low computational cost, as we detail below. In principle, our approach applies to any type of deep learning model; it also can apply to representations that are generated at several levels, as in modern transformer architectures with multiple attention layers Vaswani et al. (2017). As such it can serve as a diagnostic tool for exploring complex deep learning models.
Figure 1: Overview of our method. The SVD decomposition (I) yields concepts, whose importance (II) with respect to gender and occupation are evaluated with Sobol method. Finally, some concepts are removed, which “neutralizes” the gender information and (III) produces a fairer classifier.
2 Related work
In this section, we review pertinent literature to the two strands of our paper, bias and explainability.
2.1 Gender Bias in Natural Language Processing models
In this paper, we aim to enhance the fairness of our model by employing a prevalent approach in information theory, which involves eliminating gender-related information to prevent the model from utilizing it Kilbertus et al. (2017). This method of achieving gender fairness has its own body of literature:
A simple approach in the literature is to remove explicit gender indicators (he, she, her, etc.) directly from the dataset De-Arteaga et al. (2019). However, modern transformer models have been shown to be able to learn implicit gender indicators Devlin et al. (2018); Karita et al. (2019). This essentially limits the impact that such a naïve approach can have on this kind of models. As a matter of fact, a RoBERTa model trained without these gender indicators can still predict with upwards of 90% accuracy the gender of a person writing a short LinkedIn bio (see Appendix A).
Another approach attempts to create directly word embeddings that do not exhibit gender bias Bolukbasi et al. (2016); Garg et al. (2018); Zhao et al. (2018a, b); Caliskan et al. (2017). Although interesting in principle, Gonen and Goldberg (2019) shows these methods may camouflage gender biases in word embeddings but fail to eliminate them entirely.
In addition, gender information is usually correlated with other features and manifests itself at different levels in the sentence’s structure, effectively leaking through in the shape of implicit bias that can be easily exploited by transformer models to generate biased predictions.
A third approach, and more in line with our proposed method, consists in seeking fairness in the sense of information theory Xu et al. (2020). Specifically, we propose to use methodology from the field of explainable AI (XAI) to perform feature selection to reduce the effect of algorithmic bias Frye et al. (2020); Dorleon et al. (2022); Galhotra et al. (2022) in a human-understandable yet principled way. Indeed, our aim is to decompose the model’s embedding into an interpretable base of concepts – i.e. our features –, select those that carry the most information about gender and the least about the actual task, and suppress them from the embedding.
2.2 Explainability and Interpretability in Model Analysis
Explainability
We therefore employ explainability methods to achieve fairness in this context. Let us first provide an overview of this field.
Classical approaches analyse how each part of the input influences the model’s prediction, either through perturbation Ribeiro et al. (2016); Zeiler and Fergus (2014) or by leveraging the gradients inside the neural network Sundararajan et al. (2017). However, these methods are prone to adversarial manipulation Wang et al. (2020), perform only partial input recovery Adebayo et al. (2018), and exhibit a general lack of stability with respect to the input Ghorbani et al. (2019). Moreover, they often fail to provide full interpretability, because they provide explanations that tamper with some features but not others; in general the features tampered with do not offer a guarantee of the outcome Ignatiev et al. (2019); Marques-Silva (2023).
Rationalization approaches also play a pivotal role in enhancing our understanding Bouchacourt and Denoyer (2019); Antognini and Faltings (2021).
Asher et al. (2021) explored counterfactual approaches to isolating features that may have harmful biases. Another branch of explicability close to this one is contrastive explanations, recently used in Yin and Neubig (2022); Ferrando et al. (2023) .
The most recent approaches are concept-based methods such as TCAV Kim et al. (2018) and CRAFT Fel et al. (2023) for vision, and COCKATIEL Jourdan et al. (2023) for NLP.
Faithfulness and plausibility
When talking of explanations, we will speak of the explanans, the explaining part, and the explanandum, the part to be explained. Discussions of explainability have led to different interpretations of explainability and interpretability. In this section, we set out how we understand these terms. In the recent literature they have sometimes been used equivalently Jacovi and Goldberg (2020), but also sometimes distinguished Rudin (2019) using the concepts of faithfulness and plausibility. Plausibility has to do with how acceptable the purported explanation is to humans Jacovi and Goldberg (2020). Faithfulness, on the other hand, involves a causal connection (what is causally necessary and sufficient for the model to produce its prediction given a certain input); a faithful explanatory model is one that captures these causal relations in the explanations it generates.
To capture these causal relations, we will perform interventions removing the putative causal factors to see how this affects the model’s output. What intuitively needs to be done is an intervention that removes the impact of a certain token from the model’s prediction pipeline. However, simply masking a token from another is not sufficient to do this, because all tokens interact at multiple levels in modern transformer architectures. Blocking their interaction at one level does not block interactions at later or higher levels that can affect downstream predictions . Thus, trying to construct faithful models and causal connections requires some care.
3 Preliminaries and definitions
3.1 Methodology
The idea behind the method we propose for creating gender-neutral embedding can be summed up in 3 components: (i) Identification of occupation-related concepts within the latent space of an NLP model; (ii) evaluation of the significance of each concept in predicting gender; and (iii) removal of specific concepts to construct a gender-neutral embedding, followed by retraining a classifier to predict occupation based on this modified embedding. For a visual representation of the method, refer to Figure 1, which provides an explanatory diagram illustrating the different stages of the process.
To realize the proposed method, it is crucial to secure certain guarantees and engage in rigorous theoretical modelling. This ensures the identification and application of suitable, practical tools. This involves establishing clear criteria for selecting practical tools and methodologies, enabling the achievement of dependable and consistent results. These are defined in this section.
3.2 Notation
[…]
With $n$ individuals in the training set, we define $\bm{X}=(\bm{x}_{1},...,\bm{x}_{n})\in\mathcal{X}^{n}$, $\bm{Y}=(y_{1},...,y_{n})\in\mathcal{Y}^{n}$, and $\bm{G}=(g_{1},...,g_{n})\in\{\text{Male, Female}\}^{n}$.
The model $\bm{f}$ is conceptualized as a composition of two functions. First, $h$ encompasses all layers from the input $\bm{x}$ to a latent space, which is a transformed representation of the input data. Then, the classification layers $c$ classify these transformed data. The model can be summarized as:
$\bm{f}(\bm{x})=c\circ h(\bm{x})\,,$ (1)
where $h(\bm{x})\in\mathbb{R}^{d}$. In this space, we obtain an output embedding matrix, denoted as $\bm{A}=h(\bm{X})\in\mathbb{R}^{n\times d}$, which is also referred to as the CLS embedding in the context of an encoder transformer for classification task.
3.3 Causal underpinnings of the method
A faithful reconstruction of $\bm{f}$ should support a causal relation that can be formulated as follows: $A$ and $B$ and had $A$ not been the case, $B$ would not have been the case either Lewis (1973); Pearl (2009); Jacovi and Goldberg (2020). $A$ describes the causally necessary and “other things being equal” sufficient condition for $B$. Faithfulness implies that a particular type of input is causally necessary for the prediction at a given data point $\bm{x}$. Such a condition could take the form of one or more variables having certain values or set of variables. It can also take the form of a logical statement at a higher degree of abstraction.
A counterfactual theory is simply a collection of such statements, and for every deep learning model, there is a counterfactual theory that describes it Jacovi et al. (2021); Asher et al. (2022); Yin and Neubig (2022).
[…]
For instance, we can alternatively choose (i) $\mathscr{W}=\mathcal{X}^{n}$ the model’s input entries, (ii) $\mathscr{W}=$ the set of possible attention weights in $\bm{f}$ (assuming it’s a transformer model), (iii) $\mathscr{W}=$ the possible parameter settings for a specific intermediate layer of $\bm{f}$ –e.g., the attribution matrix as in Fel et al. (2023), (iv) $\mathscr{W}=$ the possible parameter settings for the final layer of an attention model as suggested in Wiegreffe and Pinter (2019), or (iv) sets of factors or dimensions of the final layer output matrix from the entire model Jourdan et al. (2023). Our explanatory model here follows the ideas of Jourdan et al. (2023), though it differs importantly in details. For one thing, we note that our general semantic approach allows us to have Boolean combinations even first order combinations of factors to isolate a causally necessary and sufficient explanans.
3.3.1 Counterfactual intervention for concept ranking
In part 2 of our method, we calculate the causal effect on the sensitive variable for each dimension $r$ of $\bm{U}$ and $\bm{W}$ of the $\bm{A}=\bm{U}\bm{W}$ decomposition created in part 1 of the method (more details on this decomposition in practice in section 4.1). We conduct a counterfactual intervention that involves removing the dimension or “concept” $r$. By comparing the behaviour of the model with the concept removed and the behaviour of the original model, we can assess the causal effect of the concept on the sensitive variable. As explained above, this assumes a counterfactual model where $\mathscr{W}$ corresponds to the possible variations of $\bm{A}=h(\bm{X})$ when reducing the dimensionality of its decomposition $\bm{A}=\bm{U}\bm{W}$.
In addition, this approach requires that we are really able to remove the targeted dimension, which means that the dimensions must be independent of each other.
3.4 Information theory guarantees
After performing our counterfactual intervention, we need to compute the "importance of a dimension" for the sensitive variable. Our goal is for our final embedding $\bm{A}$ to be as independent as possible from the sensitive variable $\bm{G}$.
According to information theory, if we are unable to predict a variable $\bm{G}$ from the data $\bm{A}$, it implies that this variable $\bm{G}$ does not influence the prediction of $y$ by $\bm{A}$ Xu et al. (2020). This definition is widely used in causal fairness studies Kilbertus et al. (2017), making it a logical choice for our task.
We train a classifier $c_{g}$ for the gender $\bm{G}$ prediction from the matrix $\bm{A}$ (see appendix C.2 for details on its architecture). If $c_{g}$ cannot achieve high accuracy for gender prediction, it implies that $\bm{A}$ does not contain gender-related information. The most reliable approach is thus to remove each of the dimensions found by the decomposition one by one and re-train a classifier to predict gender from this matrix. This can be computationally expensive if we have many dimensions. An alternative is to train only one classifier to predict gender on the original matrix $\bm{A}$, and then use a sensitivity analysis method on the dimensions to determine their importance for prediction. We will choose the latter for our method. However, we will still demonstrate the relevance of this approach by training a classifier each time we remove a dimension and show its effect in Section 5.2.
4 Methodology
Now that we’ve established the constraints on how to implement our method, we define the tools used for each part of the method: (i) For the decomposition that uncovers our concepts, we use the SVD decomposition. (ii) For the gender importance calculation, we train a classifier for this task and then use Sobol importance. (iii) To remove concepts, we simply remove the columns corresponding to the removed concept.
4.1 Generating concepts used for occupation classification
Firstly, we look into latent space before the last layer of the model, where the decisions are linear. We have an output embedding matrix (also called CLS embedding) of the transformer model in this space. The aim here is to employ decomposition techniques to extract both a concept coefficients matrix and a concept base matrix from latent vectors.
To illustrate this part, we discover concepts without supervision by factorizing the final embedding matrix through a Singular Value Decomposition. We choose this technique, but it is possible to do other decompositions, like PCA Wold et al. (1987), RCA Bingham and Mannila (2001) or ICA Comon (1994).
Singular Value Decomposition (SVD)
The embedding matrix $\bm{A}\in\mathbb{R}^{n\times d}$ can be decomposed using an SVD as:
$\bm{A}=U_{0}\Sigma_{0}V_{0}^{\intercal}=U_{0}W_{0}\,,$
where $U_{0}\in\mathbb{R}^{n\times d}$ contains the left singular vectors, $V_{0}\in\mathbb{R}^{d\times d}$ contains the right singular vectors, and $\Sigma_{0}\in\mathbb{R}^{d\times d}$ is a diagonal matrix containing the singular values. Note that $U_{0}$ and $V_{0}$ are orthonormal matrices, which means that $U_{0}^{\intercal}U_{0}=Id$ and $V_{0}^{\intercal}V_{0}=Id$.
An SVD can be used to explain the main sources of variability of the decomposed matrix, here $\bm{A}$. The matrix $W_{0}^{-1}=V_{0}\Sigma_{0}^{-1}$ is indeed considered as a linear projector of $\bm{A}$, and $U_{0}$ is the projection of $\bm{A}$ on a basis defined by $V_{0}$. The amount of variability, or more generally, of information that has been captured in $\bm{A}$ by a column of $U_{0}$ is directly represented by the corresponding singular value in $\Sigma$. Using only the $r$ largest singular values of $\Sigma_{0}$, with $r<<d$, therefore allows capturing as much variability in $\bm{A}$ as possible with a reduced amount of dimensions. We then decompose $\bm{A}$ as:
$\bm{A}\approx\bm{U}\bm{W}\,,$ (2)
where $\bm{W}=\Sigma V^{\intercal}$ is a compressed version of $W_{0}$, in the sense that $\Sigma\in\mathbb{R}^{r,r}$ is the diagonal matrix containing the $r$ largest singular values of $\Sigma$ and $V\in\mathbb{R}^{d,r}$ contains the $r$ corresponding right singular vectors. The matrix $\bm{U}\in\mathbb{R}^{n,r}$ consequently contains the $r$ right singular vectors which optimally represent the information contained in $\bm{A}$ according to the SVD decomposition. In this paper, we refer to these vectors as the extracted concepts of $\bm{A}$. Python implementation details are in Appendix D.
4.2 Estimating concept importance with Sobol Indices
We have seen in Eq. (1) that $c$ is used to predict $\bm{Y}$ based on $\bm{A}$, and we discussed in section 3.4, that the non-linear classifier $c_{g}$ also uses $\bm{A}$ as input but predicts the gender $\bm{G}$. We also recall that $\bm{A}$ is approximated by $\bm{U}\bm{W}$ in Eq. (2), and that the columns of $\bm{U}$ represent the so-called concepts extracted from $\bm{A}$.
We then evaluate the concepts’ importance by applying a feature importance technique on $\bm{G}$ using $c_{g}$, and on $\bm{Y}$ using $c$. More specifically, Sobol indices Sobol (1993) are used to estimate the concepts’ importance, as in Jourdan et al. (2023). A key novelty of the methodology developed in our paper is to simultaneously measure each concept importance for the outputs $\bm{G}$ and $\bm{Y}$, and not only $\bm{Y}$ as proposed in Jourdan et al. (2023).
Figure 2: Co-importance plot for $r=20$ dimensions with respect to Occupation and Gender labels, computed with the Sobol method. The colour is based on the angle $a=90-\frac{2}{\pi}\arctan{(\frac{y}{x})}$. The angles with extreme values correspond to concepts with high importance for gender, but comparatively low importance for occupation. Their removal yields a favourable tradeoff in accuracy/fairness.
To estimate the importance of a concept $\bm{U}_{i}$ for gender (resp. for the label), we measure the fluctuations of the model output $c_{g}(\bm{U}\bm{W})$ (resp. $c(\bm{U}\bm{W})$) in response to perturbations of concept coefficient $\bm{U}_{i}$
We then propagate this perturbed activation to the model output $\bm{G}=c_{g}(\tilde{\bm{A}})$ (resp. $\bm{Y}=c(\tilde{\bm{A}})$). We can capture the importance that a concept might have as a main effect – along with its interactions with other concepts – on the model’s output by calculating the expected variance that would remain if all the concepts except the i were to be fixed. This yields the general definition of the total Sobol indices. An important concept will have a large variance in the model’s output, while an unused concept will barely change it.
We explain the total Sobol indices for each concept in details in Appendix E, along with our perturbation strategy.
4.3 Neutralizing the embedding with importance-based concept removal
In this final section, we propose to eliminate $k$ concepts that exhibit significant gender dependence within the final embedding matrix, with the objective of improving its gender-neutrality. This provides us with an improved embedding which can serve as a base to learn a fairer classifier to predict downstream tasks, and thus, a more gender-neutral model overall.
This concept-removal procedure consists of choosing the $k$ concepts that maximize the ratio between the importance for gender prediction and for the task. This graphically corresponds to searching for the concepts whose angle with respect to the line of equal importance for gender and task prediction are the highest (see Fig. 2). This strategy allows us to work on a trade-off between the model’s performance and its fairness.
In practice, we sort the concepts $\bm{U}_{1}$,$\bm{U}_{2}$,….$\bm{U}_{r}$ according to the second strategy in descending order and we delete the $k$ first concepts. This essentially leaves us with a concept base $\tilde{\bm{U}}=(\bm{U}_{k+1},...,\bm{U}_{r})$. This same transformation can be applied to $\bm{W}$’s lines, yielding $\tilde{\bm{W}}$. From the matrix obtained from the matrix product $\tilde{\bm{U}}\tilde{\bm{W}}$, we retrain a classifier $\tilde{c}$ to predict $\bm{Y}$ (refer to Appendix C.2 for more details).
Remark:
The more concepts we remove, the less we will be able to predict the gender, so the fairer the model will be. However, this also entails a decrease on the amount of information that is encoded in the embedding and that can be used for downstream tasks. The parameter $k$ must therefore be chosen to comply with our specific application’s constraints.
Figure 3: Accuracy drop when $k$ concepts are removed for an SVD with $r=20$ dimensions. In Figure 2(a), (Top) the occupation accuracy, (Bottom) the gender accuracy. Figure 2(b) shows occupation accuracy as a function of gender accuracy drop. The gender accuracy drop is reported relatively to the Baseline1, the classifier trained on Bios with all the concepts.
5 Results
5.1 Dataset and models
To illustrate our method, we use the Bios dataset De-Arteaga et al. (2019), which contains about 440K biographies (textual data) with labels for the genders (M or F) of the authors, and their occupations (28 classes). From the Bios dataset, we also create a second dataset Bios-neutral by removing all the explicit gender indicators. Bios-neutral will be used as a baseline to which we can compare our method. We leave the details of how this dataset preprocessing was performed in Appendix B.
For both datasets, we train a RoBERTa (Liu et al., 2019a) model to predict the occupation task. We leave the details on how these models were trained in Appendix C.
5.2 Results analysis - Quantitative part
Figure 4: Understanding concepts extracted from the model trained on Bios. Concept 1 prominently features explicit gender indicators like ’she’ and ’his’, crucial for gender identification but irrelevant for occupation. Concept 8 integrates information about gender-unbalanced professions, aiding in predicting both, an individual’s gender and occupation. Concept 19 focuses on information related to the professor occupation, a gender-balanced class and most represented in the dataset.
Initially, we will examine the progression of our model’s performance for each dimension removed (one by one) utilizing our methodology.
Following the precepts of information theory, we will assess the effectiveness of our bias mitigation strategy by measuring the model’s capacity to predict the gender and the task as we progressively remove concepts.
We showcase in Fig. 3 the outcome of removing the $k$ most important concepts for $k\in\llbracket 0,19\rrbracket$ (we have $r=20$ dimensions) in terms of different metrics: accuracy on gender prediction and on predicting occupation (downstream task). We compare our technique to two baselines: Baseline1 – RoBERTa model trained on Bios dataset (i.e. when $k=0$) –, and Baseline2 – RoBERTa model trained on Bios-neutral dataset.
From Figure 3, simply by removing one concept, we achieve significant gains in the gender neutrality of our model, as it predicts gender with much lower accuracy (82.3% compared to the initial 96%) while maintaining its performance in occupation prediction (accuracy decreasing from 86.4% to 86.3%). Starting from the removal of the fourth dimension, the accuracy of gender prediction falls below the threshold of Baseline2, with 79.4% accuracy compared to 79.8% for Baseline2, while maintaining an accuracy of 86.3% for occupation (compared to 86% for Baseline2). Moreover, our method, in addition to being explainable, time-efficient, and cost-effective, exhibits greater effectiveness.
After removing these first four concepts, we enter a regime where removing concepts continues to gain us fairness (as the accuracy of the gender task continues to fall), while the accuracy of the occupation remains unchanged. This regime is therefore the one with the best trade-off between accuracy and fairness (shown in green here). From concept 12 onwards, the accuracy of occupation begins to decline, as we look more and more at concepts with a high importance for occupation (according to Sobol), and we enter a new regime where we gain more and more in fairness, but where the accuracy of our task collapses. It may be worthwhile deleting a fairly large number of concepts to get into this zone, but only if we are willing to pay the price in terms of prediction quality.
It is important to note that in this specific application, the optimal fairness-accuracy trade-off corresponds to accuracy on gender prediction of 62% as that’s the accuracy one gets when learning a model to predict the gender with the occupation labels as only information.
5.3 Concepts removed analysis - Explainability part
In this section, we interpret the dimensions of the Singular Value Decomposition (SVD) as valuable concepts for predicting our two tasks: gender and occupation. For this, we use the interpretability part of the COCKATIEL Jourdan et al. (2023) method on concepts/dimensions previously found with the SVD and ranked with the Sobol method by importance for gender and occupation.
Concept interpretability
COCKATIEL modifies the Occlusion method Zeiler and Fergus (2014), which operates by masking each word and subsequently observing the resultant impact on the model’s output. In this instance, to infer the significance of each word pertaining to a specific concept, words within a sentence are obscured, and the influence of the modified sentence (devoid of the words) on the concept is measured. This procedure can be executed at either the word or clause level, meaning it can obscure individual words or entire clauses, yielding explanations of varying granularity contingent upon the application.
Concepts interpretation
For the dimensions of the SVD to be meaningful for explanation, they must first be meaningful for the model’s prediction, whether it be the model predicting gender or the one predicting occupation. We rank the concepts w.r.t their removal order, and we focus on three important ones:
Concept 1, with a gender importance of 32%, and an occupation importance of 5%; Concept 8, with a gender importance of 16%, and an occupation importance of 12%; Concept 19, with a gender importance of 12%, and an occupation importance of 33%.
These three concepts are crucial for the prediction of at least one of our two tasks. Moreover, they are representative of our problem statement as Concept 1 is more important for gender, Concept 8 is as important to gender as occupation, and Concept 19 is more important for occupation. We illustrate several examples of explanations for these three concepts in Figure 4 to enable qualitative analysis.
According to the importance per dimension calculated in part 2 of our method, concept 1 is almost exclusively predictive of gender. The most significant words for the concept of this dimension are explicit indicators of traditional gender.
Concept 8 is capable of predicting both gender and occupation. Within its concept, we find examples of biographies predominantly related to unbalanced professions such as dietitian (92.8% women), chiropractor (26.8% women), software engineer (16.2% women), and model (83% women). It is logical that the concept allowing the prediction of these professions also enables the prediction of gender, as it is strongly correlated in these particular professions.
Concept 19, which strongly predicts occupation, focuses on biographies predicting professors (constituting more than 29.7% of all individuals). This class is gender-balanced (with 45.1% women), providing little information on gender.
6 Conclusion
In this paper, we introduced a method for neutralizing the presence of bias in transformer model’s embeddings in an interpretable and cost-efficient way through concept-based explainability. In this study, we present a method applied to gender-related information, yet inherently adaptable to other sensitive variables. Indeed, by removing concepts that are mostly influential for gender prediction, we obtain embeddings that rely considerably less on gender information for generating predictions, and at the same time, remain highly accurate at the task. Furthermore, by doing so via explainable AI, we perform interventions that can be easily interpreted by humans, and thus gain trust in the bias mitigation process and insights into what produced the bias in the first place. Finally, thanks to its low computation cost, it can be easily applied to pre-trained, fine-tuned and even embeddings accessed through inference APIs, thus allowing our method to be applied to any model in a post-hoc manner.
Limitations
Herein, we propose a novel approach that serves as a framework that can be applied to various concept extraction techniques and feature importance computation methods. This means that the SVD could be exchanged for other dimensionality reduction techniques (and thus, change the properties of the concept base and coefficients), or Shapley values could be preferred over Sobol indices for the relaxation on the independence assumption, for istance. In addition, our application was gender bias mitigation, but it could be applied to any other type of bias as long as the sensitive variable is available or can be estimated.
Acknowledgements
We thank the ANR-3IA Artificial and Natural Intelligence Toulouse Institute (ANITI) funded by the ANR-19-PI3A-0004 grant for research support. This work was conducted as part of the DEEL [^1] https://www.deel.ai project.
Appendix A Gender prediction with RoBERTa
In this section, we study the impact on gender prediction of removing explicit gender indicators with a transformer model. To do this, we train a model on a dataset and we train a model with the same architecture on modified version of the dataset where we remove the explicit gender indicators. We train these models to predict the gender and we compare both of these model’s accuracies.
In our example, we use the Bios dataset De-Arteaga et al. (2019), which contains about 400K biographies (textual data). For each biography, we have the gender (M or F) associated. We use each biography to predict the gender, and then we clean the dataset and we create a second dataset without the explicit gender indicators (following the protocol in Appendix B).
We train a RoBERTa model, as explained in Appendix C.1. Instead of training the model on the occupations, we place ourselves directly in the case where we train the RoBERTa to predict gender (with the same parameters). We train both models over 5 epochs.
In figure 5, we go from 99% accuracy for the baseline to 90% accuracy for the model trained on the dataset without explicit gender indicators. Even with a dataset without explicit gender indicators, a model like RoBERTa performs well on the gender prediction task.
Figure 5: Convergence curves (accuracy) during training for the gender classification task, 5 epochs. (On the left) RoBERTa model trained on Bios; (On the right): RoBERTa model trained on Bios without explicit gender indicators.
Appendix B Dataset pre-processing protocol
We use Bios dataset De-Arteaga et al. (2019), which contains about 400K biographies (textual data). For each biography, Bios specifies the gender (M or F) of its author as well as its occupation (among a total of 28 possible occupations, categorical data).
As shown in Figure 6, this dataset contains heterogeneously represented occupations. Although the representation of some occupations is relatively well-balanced between males and females, other occupations are particularly unbalanced (see Figure 7).
Figure 6: Number of biographies for each occupation by gender on the total Bios dataset De-Arteaga et al. (2019).
Figure 7: Co-occurrence of occupation and gender for Bios dataset De-Arteaga et al. (2019). Solely predicting the gender from the occupation yields a minimum of 62% accuracy, which is the minimum baseline for an accurate but unbiased model. Any classifier with lower accuracy on gender prediction must have accuracy on occupation prediction lower than $100\%$.
Cleaning dataset
To clean the dataset, we take all biographies and apply these modifications:
  • we remove mail and URL,
  • we remove dot after pronouns and acronyms,
  • we remove double "?", "!" and ".",
  • we cut all the biographies at 512 tokens by checking that they are cut at the end of a sentence and not in the middle.
Creating the Bios-neutral dataset
From Bios, we create a new dataset without explicit gender indicators, called Bios-neutral.
At first, we tokenize the dataset, then:
  • If the token is a first name, we replace it with "Sam" (a common neutral first name).
  • If the token is a dictionary key (on the dictionary defined below), we replace it with its values.
To determine whether the token is a first name, we use the list of first names: usna.edu
The dictionary used in the second part is based on the one created in Field and Tsvetkov (2020).
Appendix C Models details and convergence curves
C.1 RoBERTa training
RoBERTa base training
We use a RoBERTa model (Liu et al., 2019a), which is based on the transformers architecture and is pretrained with the Masked language modelling (MLM) objective. We specifically used a RoBERTa base model pretrained by Hugging Face. All information related to how it was trained can be found in Liu et al. (2019b). It can be remarked, that a very large training dataset was used to pretrain the model, as it was composed of five datasets: BookCorpus Zhu et al. (2015), a dataset containing 11,038 unpublished books; English Wikipedia (excluding lists, tables and headers); CC-News Mackenzie et al. (2020) which contains 63 millions English news articles crawled between September 2016 and February 2019; OpenWebText Radford et al. (2019) an open-source recreation of the WebText dataset used to train GPT-2; Stories Trinh and Le (2018) a dataset containing a subset of CommonCrawl data filtered to match the story-like style of Winograd schemas. Pre-training was performed on these data by randomly masking 15% of the words in each of the input sentences and then trying to predict the masked words.
Figure 8: Convergence curves during training for occupation classification task, 15 epochs. Baseline1: RoBERTa model trained on Bios; Baseline2: RoBERTa model trained on Bios without explicit gender indicators. (On the left) accuracy curves; (on the right) loss curves.
Occupation prediction task
After pre-training RoBERTa parameters on this huge dataset, we then trained it on the 400.000 biographies of the Bios dataset. The training was performed with PyTorch on 2 GPUs (Nvidia Quadro RTX6000 24GB RAM) for 10 epochs (we train 15 epochs but we stop at 10 epochs) with a batch size of 8 observations and a sequence length of 512 words. The optimizer was Adam with a learning rate of 1e-6, $\beta_{1}=0.9$, $\beta_{2}=0.98$, and $\epsilon=1e6$. We split dataset into 70% for training, 10% for validation, and 20% for testing.
The convergence curves of the model are in Figure 8.
C.2 Classifier heads training
The importance estimation in part 2 of our method requires a classifier to be trained on top of the features. An ideal classifier will leverage the maximum amount possible of information about gender or occupation that can be realistically extracted from latent features, in the spirit of the maximum information that can be extracted under computational constraints Xu et al. (2020). Hence, it is crucial to optimize an expressive model to the highest possible performance.
In the latent space of the feature extractor, the decisions are usually computed with a linear classifier. However, when operating on latent space with removed concepts or when working with a task on which the transformer head has not been fine-tuned, nothing guarantees that the task can be solved with linear probes. Hence, we chose a more expressive model, namely a two-layer perceptron with ReLU non-linearities, and architecture $768\rightarrow 128\rightarrow C$ with $C$ the number of classes. The network is trained by minimizing the categorical cross-entropy on the train set, with Softmax activation to convert the logits into probabilities. We use Adam with a learning rate chosen between $10^{-5}$ and $10^{-1}$ with cross-validation on the validation set. Note that the paper reports the accuracy on the test set.
The gender classifier $c_{g}$ and the occupation classifer $\tilde{c}$ are both trained with this same protocol.
Appendix D Implementation details for SVD
We used the SVD implementation in sparse.linalg.svds provided in the Scipy package. This function makes use of an incremental strategy to only estimate the first singular vectors of $\bm{A}$, making the truncated SVD scalable to large matrices. Note that this implementation turned out to obtain the most robust decompositions of $\bm{A}$ after an empirical comparison between torch.svd, torch.linalg.svd, scipy.linalg.svd, and sklearn.decomposition.TruncatedSVD on our data.
Appendix E Sobol Indices - details
In this section, we define the classic Total Sobol indices and how they are calculated using our method. In practice, these indices can be calculated very efficiently Marrel et al. (2009); Saltelli et al. (2010); Janon et al. (2014) with Quasi-Monte Carlo sampling and the estimator explained below.
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space of possible concept perturbations. To build these concept perturbations, we use $\boldsymbol{M}=\left(M_{1},\ldots,M_{r}\right)\in\mathcal{M}\subseteq[0,1]^{r}$, i.i.d. stochastic masks where $M_{i}\sim\mathcal{U}(\left[0,1\right]),\forall i\in 1,...,r$. We define concept perturbation $\tilde{\bm{U}}=\boldsymbol{\pi}(\bm{U},\boldsymbol{M})$ with the perturbation operator $\boldsymbol{\pi}(\bm{U},\boldsymbol{M})=\bm{U}\odot\boldsymbol{M}$ with $\odot$ the Hadamard product (that takes in two matrices of the same dimensions and returns a matrix of the multiplied corresponding elements).
We denote the set $\mathcal{U}=\{1,\ldots,r\}$, $\boldsymbol{u}$ a subset of $\mathcal{U}$, its complementary $\sim\boldsymbol{u}$ and $\mathbb{E}(\cdot)$ the expectation over the perturbation space. We define $\phi:\mathcal{A}\rightarrow\mathbb{R}$, a function that takes the perturbations $\bm{U}$ from the last layer and applies the difference between the model’s two largest output logits – i.e. $\phi(\bm{U})=l_{(1)}(c(\bm{U}))-l_{(2)}(c(\bm{U}))$ where $c$ is the classification head of the model (when we do Sobol for the occupation task, it is $c$, and when we do Sobol for the gender task, it is $c_{g}$) and $l_{(1)}$ and $l_{(2)}$ represent the highest and second highest logit values respectively. We assume that $\phi\in\mathbb{L}^{2}(\mathcal{A},\mathbb{P})$ – i.e. $|\mathbb{E}(\phi(\boldsymbol{U}))|<+\infty$.
The Hoeffding decomposition provides $\phi$ as a function of summands of increasing dimension, denoting $\phi_{\boldsymbol{u}}$ the partial contribution of the concepts $\boldsymbol{U}_{\boldsymbol{u}}=\left(U_{i}\right)_{i\in\boldsymbol{u}}$ to the score $\phi(\boldsymbol{U})$:
$\phi(\boldsymbol{U})$ $=\phi_{\varnothing}$ (3)
$+\sum_{i=1}^{r}\phi_{i}\left(U_{i}\right)$
$+\sum_{1\leqslant i<j\leqslant r}\phi_{i,j}\left(U_{i},U_{j}\right)+\cdots$
$+\phi_{1,\ldots,r}\left(U_{1},\ldots,U_{r}\right)$
$=\sum_{\boldsymbol{u}\subseteq\mathcal{U}}\phi_{\boldsymbol{u}}\left(\boldsymbol{U}_{\boldsymbol{u}}\right)$
Eq. 3 consists of $2^{r}$ terms and is unique under the orthogonality constraint:
$\quad\mathbb{E}\left(\phi_{\boldsymbol{u}}\left(\boldsymbol{U}_{\boldsymbol{u}}\right)\phi_{\boldsymbol{v}}\left(\boldsymbol{U}_{\boldsymbol{v}}\right)\right)=0,$
$\forall(\boldsymbol{u},\boldsymbol{v})\subseteq\mathcal{U}^{2}\text{ s.t. }\boldsymbol{u}\neq\boldsymbol{v}$
Moreover, thanks to orthogonality, we have $\phi_{\boldsymbol{u}}\left(\boldsymbol{U}_{\boldsymbol{u}}\right)=\mathbb{E}\left(\phi(\boldsymbol{U})\mid\boldsymbol{U}_{\boldsymbol{u}}\right)-\sum_{\boldsymbol{v}\subset\boldsymbol{u}}\phi_{\boldsymbol{v}}\left(\boldsymbol{U}_{\boldsymbol{v}}\right)$ and we can write the model variance as:
$\mathbb{V}(\phi(\boldsymbol{U}))$ $=\sum_{i}^{r}\mathbb{V}\left(\phi_{i}\left(U_{i}\right)\right)$ (4)
$+\sum_{1\leqslant i<j\leqslant r}\mathbb{V}\left(\phi_{i,j}\left(U_{i},U_{j}\right)\right)$
$+\ldots+\mathbb{V}\left(\phi_{1,\ldots,r}\left(U_{1},\ldots,U_{r}\right)\right)$
$=\sum_{\boldsymbol{u}\subseteq\mathcal{U}}\mathbb{V}\left(\phi_{\boldsymbol{u}}\left(\boldsymbol{U}_{\boldsymbol{u}}\right)\right)$
Eq. 4 allows us to write the influence of any subset of concepts $\boldsymbol{u}$ as its own variance. This yields, after normalization by $\mathbb{V}(\phi(\boldsymbol{U}))$, the general definition of Sobol’ indices.
Definition
Sobol’ indices Sobol (1993). The sensitivity index $\mathcal{S}_{\boldsymbol{u}}$ which measures the contribution of the concept set $\boldsymbol{U}_{\boldsymbol{u}}$ to the model response $f(\boldsymbol{U})$ in terms of fluctuation is given by:
$\mathcal{S}_{\boldsymbol{u}}=\frac{\mathbb{V}\left(\phi_{\boldsymbol{u}}\left(\boldsymbol{U}_{\boldsymbol{u}}\right)\right)}{\mathbb{V}(\phi(\boldsymbol{U}))}=\$ (5)
$\frac{\mathbb{V}\left(\mathbb{E}\left(\phi(\boldsymbol{U})\mid\boldsymbol{U}_{\boldsymbol{u}}\right)\right)-\sum_{\boldsymbol{v}\subset\boldsymbol{u}}\mathbb{V}\left(\mathbb{E}\left(\phi(\boldsymbol{U})\mid\boldsymbol{U}_{\boldsymbol{v}}\right)\right)}{\mathbb{V}(\phi(\boldsymbol{U}))}$
Sobol’ indices provide a numerical assessment of the importance of various subsets of concepts in relation to the model’s decision-making process. Thus, we have: $\sum_{\boldsymbol{u}\subseteq\mathcal{U}}\mathcal{S}_{\boldsymbol{u}}=1$.
Additionally, the use of Sobol’ indices allows for the efficient identification of higher-order interactions between features. Thus, we can define the Total Sobol indices as the sum of of all the Sobol indices containing the concept $i:\mathcal{S}_{T_{i}}=\sum_{\boldsymbol{u}\subseteq\mathcal{U},i\in\boldsymbol{u}}\mathcal{S}_{\boldsymbol{u}}$. So we can write:
Definition
Total Sobol indices. The total Sobol index $\mathcal{S}{T_{i}}$, which measures the contribution of a concept $\bm{U}_{i}$ as well as its interactions of any order with any other concepts to the model output variance, is given by:
$\mathcal{S}_{T_{i}}$ $=\frac{\mathbb{E}_{\bm{M}_{\sim i}}(\mathbb{V}_{M_{i}}(\bm{Y}|\bm{M}_{\sim i}))}{\mathbb{V}(\bm{Y})}$ (6)
$=\frac{\mathbb{E}_{\bm{M}_{\sim i}}(\mathbb{V}_{M_{i}}(\phi((\bm{U}\odot\bm{M})\bm{W})|\bm{M}_{\sim i}))}{\mathbb{V}(\phi((\bm{U}\odot\bm{M})\bm{W}))}.$ (7)
In practice, our implementation of this method remains close to what is done in COCKATIEL Jourdan et al. (2023), with the exception of changes to the $\phi$ function created here to manage concepts found in all the classes at the same time (unlike COCKATIEL, which concentrated on finding concepts in a per-class basis).

### v4

@@
-TaCo: Targeted Concept Removal in Output Embeddings for NLP via Information Theory and Explainability
+TaCo: Targeted Concept Erasure Prevents Non-Linear Classifiers From Detecting Protected Attributes
 Abstract
-The fairness of Natural Language Processing (NLP) models has emerged as a crucial concern.
-Information theory indicates that to achieve fairness, a model should not be able to predict sensitive variables, such as gender, ethnicity, and age.
-However, information related to these variables often appears implicitly in language, posing a challenge in identifying and mitigating biases effectively.
-To tackle this issue, we present a novel approach that operates at the embedding level of an NLP model, independent of the specific architecture.
-Our method leverages insights from recent advances in XAI techniques and employs an embedding transformation to eliminate implicit information from a selected variable.
-By directly manipulating the embeddings in the final layer, our approach enables a seamless integration into existing models without requiring significant modifications or retraining.
-In evaluation, we show that the proposed post-hoc approach significantly reduces gender-related associations in NLP models while preserving the overall performance and functionality of the models.
-An implementation of our method is available: https://github.com/fanny-jourdan/TaCo
+Ensuring fairness in NLP models is crucial, as they often encode sensitive attributes like gender and ethnicity, leading to biased outcomes.
+Current concept erasure methods attempt to mitigate this by modifying final latent representations to remove sensitive information without retraining the entire model.
+However, these methods typically rely on linear classifiers, which leave models vulnerable to non-linear adversaries capable of recovering sensitive information.
+We introduce Targeted Concept Erasure (TaCo), a novel approach that removes sensitive information from final latent representations, ensuring fairness even against non-linear classifiers.
+Our experiments show that TaCo outperforms state-of-the-art methods, achieving greater reductions in the prediction accuracy of sensitive attributes by non-linear classifier while preserving overall task performance.
+Code is available on GitHub.
 1 Introduction
-The multiple and impressive successes of Artificial Intelligence and Machine Learning have made them more and more ubiquitous in applications where a model’s decisions has genuine repercussions on people’s lives.
-These decisions, however, typically come with biases that can lead to unfair, unethical consequences.
-To guard against such consequences, we must identify and understand the biased information these models utilize.
-This links the task of ensuring an ethical comportment of a model with its interpretability or explainability: if we can show the real reasons or features why a model makes the predictions or recommendations it does, then we can control for biases among these features.
-In this paper, we propose a method for finding factors that influence model predictions using an algebraic decomposition of the latent representations of the model into orthogonal dimensions.
-Using the method of Singular Value Decomposition (SVD), which is well suited to our particular test case, we isolate various dimensions that contribute to a prediction.
-As our case study, we study the use of gender information in predicting occupations in the Bios dataset De-Arteaga et al.
-(2019) by a RoBERTa model Liu et al.
-(2019a).
-By intervening on the latent representations to remove certain dimensions and running the model on the edited representations, we show first a causal connection between various dimensions and model predictions about gender and separately about occupation.
-This enables us then to eliminate dimensions in the decomposition that convey gender bias in the decision without much affecting the model’s predictive performance on occupations.
-Because SVD gives us orthogonal dimensions, we know that once we have removed a dimension, the information it contains is no longer in the model.
-Crucially, we show how to link various dimensions in the SVD decomposition of the latent representation with concepts like gender and its linguistic manifestations that matter to fairness, providing not only a faithful explanation of the model’s behaviour but a humanly understandable or plausible one as well.
-Faithfulness provides a causal connection and produces factors necessary and sufficient for the model to produce its prediction given a certain input.
-Plausibility has to do with how acceptable the purported explanation is to humans Jacovi and Goldberg (2020).
-In addition, we demonstrate that our method can reach close to the theoretically ideal point of minimizing algorithmic bias while preserving strong predictive accuracy, and we offer a way of quantifying various operations on representations.
-Our method also has a relatively low computational cost, as we detail below.
-In principle, our approach applies to any type of deep learning model; it also can apply to representations that are generated at several levels, as in modern transformer architectures with multiple attention layers Vaswani et al.
-(2017).
-As such it can serve as a diagnostic tool for exploring complex deep learning models.
-Figure 1: Overview of our method.
-The SVD decomposition (I) yields concepts, whose importance (II) with respect to gender and occupation are evaluated with Sobol method.
-Finally, some concepts are removed, which “neutralizes” the gender information and (III) produces a fairer classifier.
+Artificial Intelligence models are increasingly deployed in critical domains but often unintentionally encode sensitive attributes such as gender, ethnicity, and age, leading to unfair and unethical outcomes.
+Ensuring fairness in these models is imperative; and according to information theory, a fair model should be incapable of predicting these sensitive variables, regardless of whether the predictor is linear or non-linear.
+Concept erasure methods have been developed to meet this desideratum.
+These methods modify the representation in the latent space of NLP models just before classification, attempting to remove information about the sensitive variable in this representation, to enable fairer classification without the need to re-train the entire model.
+These methods aim to hide sensitive information in final latent space of models, but generally evaluate their effectiveness using linear classifiers such as logistic regression (Ravfogel et al., 2020, 2022; Belrose et al., 2023).
+While these approaches reduce the ability of linear models to recover sensitive information, they often fail to eliminate it entirely, leaving it accessible to non-linear models (Ravfogel et al., 2023).
+This is a significant limitation for many concept erasure methods to achieve true fairness.
+To address this limitation, we propose Targeted Concept Erasure (TaCo), a novel concept erasure method that effectively removes sensitive information from final latent representations, ensuring fairness even against non-linear adversaries.
+Our approach involves three key steps:
+(i) Concept Discovery: We identify relevant concepts within the latent space that are associated with the sensitive attribute.
+(ii) Concept Ranking: We rank these concepts based on their importance to both the sensitive variable and the target label.
+This ranking allows us to find the optimal trade-off between fairness and task performance.
+(iii) Concept Erasure: We systematically remove concepts, prioritizing those that are most influential to the sensitive attribute but have minimal impact on the primary task prediction.
+We evaluate our method on the Bios dataset (De-Arteaga et al., 2019) using a variety of pre-trained language models, including RoBERTa (Liu et al., 2019a), DistilBERT (Sanh et al., 2019), DeBERTa (He et al., 2020), and T5 (Raffel et al., 2020).
+Our experiments demonstrate that TaCo significantly outperforms state-of-the-art concept erasure methods in preventing non-linear classifiers, such as multilayer perceptrons (MLPs), from predicting sensitive attributes.
+Specifically, TaCo achieves a greater reduction in the accuracy of predicting sensitive variables across all tested models, without compromising overall (intended) task performance.
+In addition to enhancing fairness, our method offers interpretability benefits.
+By identifying and manipulating specific concepts, we answer the question: "Which concepts have an influence?".
+And we further extend our approach by applying concept-based explainability techniques –specifically, COCKATIEL (Jourdan et al., 2023)– directly on the identified concepts.
+This allows us to also answer the question: "Why did this concept have an influence?", providing deeper understanding and transparency into the model’s decision-making process.
 2 Related work
-In this section, we review pertinent literature to the two strands of our paper, bias and explainability.
-2.1 Gender Bias in Natural Language Processing models
-In this paper, we aim to enhance the fairness of our model by employing a prevalent approach in information theory, which involves eliminating gender-related information to prevent the model from utilizing it Kilbertus et al.
-(2017).
-This method of achieving gender fairness has its own body of literature:
-A simple approach in the literature is to remove explicit gender indicators (he, she, her, etc.) directly from the dataset De-Arteaga et al.
-(2019).
-However, modern transformer models have been shown to be able to learn implicit gender indicators Devlin et al.
-(2018); Karita et al.
-(2019).
-This essentially limits the impact that such a naïve approach can have on this kind of models.
-As a matter of fact, a RoBERTa model trained without these gender indicators can still predict with upwards of 90% accuracy the gender of a person writing a short LinkedIn bio (see Appendix A).
-Another approach attempts to create directly word embeddings that do not exhibit gender bias Bolukbasi et al.
-(2016); Garg et al.
-(2018); Zhao et al.
-(2018a, b); Caliskan et al.
-(2017).
-Although interesting in principle, Gonen and Goldberg (2019) shows these methods may camouflage gender biases in word embeddings but fail to eliminate them entirely.
-In addition, gender information is usually correlated with other features and manifests itself at different levels in the sentence’s structure, effectively leaking through in the shape of implicit bias that can be easily exploited by transformer models to generate biased predictions.
-A third approach, and more in line with our proposed method, consists in seeking fairness in the sense of information theory Xu et al.
-(2020).
-Specifically, we propose to use methodology from the field of explainable AI (XAI) to perform feature selection to reduce the effect of algorithmic bias Frye et al.
-(2020); Dorleon et al.
-(2022); Galhotra et al.
-(2022) in a human-understandable yet principled way.
-Indeed, our aim is to decompose the model’s embedding into an interpretable base of concepts – i.e.
-our features –, select those that carry the most information about gender and the least about the actual task, and suppress them from the embedding.
-2.2 Explainability and Interpretability in Model Analysis
-Explainability
-We therefore employ explainability methods to achieve fairness in this context. Let us first provide an overview of this field.
-Classical approaches analyse how each part of the input influences the model’s prediction, either through perturbation Ribeiro et al.
-(2016); Zeiler and Fergus (2014) or by leveraging the gradients inside the neural network Sundararajan et al.
-(2017).
-However, these methods are prone to adversarial manipulation Wang et al.
-(2020), perform only partial input recovery Adebayo et al.
-(2018), and exhibit a general lack of stability with respect to the input Ghorbani et al.
-(2019).
-Moreover, they often fail to provide full interpretability, because they provide explanations that tamper with some features but not others; in general the features tampered with do not offer a guarantee of the outcome Ignatiev et al.
-(2019); Marques-Silva (2023).
-Rationalization approaches also play a pivotal role in enhancing our understanding Bouchacourt and Denoyer (2019); Antognini and Faltings (2021).
-Asher et al.
-(2021) explored counterfactual approaches to isolating features that may have harmful biases.
-Another branch of explicability close to this one is contrastive explanations, recently used in Yin and Neubig (2022); Ferrando et al.
-(2023) .
-The most recent approaches are concept-based methods such as TCAV Kim et al. (2018) and CRAFT Fel et al. (2023) for vision, and COCKATIEL Jourdan et al. (2023) for NLP.
-Faithfulness and plausibility
-When talking of explanations, we will speak of the explanans, the explaining part, and the explanandum, the part to be explained.
-Discussions of explainability have led to different interpretations of explainability and interpretability.
-In this section, we set out how we understand these terms.
-In the recent literature they have sometimes been used equivalently Jacovi and Goldberg (2020), but also sometimes distinguished Rudin (2019) using the concepts of faithfulness and plausibility.
-Plausibility has to do with how acceptable the purported explanation is to humans Jacovi and Goldberg (2020).
-Faithfulness, on the other hand, involves a causal connection (what is causally necessary and sufficient for the model to produce its prediction given a certain input); a faithful explanatory model is one that captures these causal relations in the explanations it generates.
-To capture these causal relations, we will perform interventions removing the putative causal factors to see how this affects the model’s output.
-What intuitively needs to be done is an intervention that removes the impact of a certain token from the model’s prediction pipeline.
-However, simply masking a token from another is not sufficient to do this, because all tokens interact at multiple levels in modern transformer architectures.
-Blocking their interaction at one level does not block interactions at later or higher levels that can affect downstream predictions .
-Thus, trying to construct faithful models and causal connections requires some care.
+In information theory, a common approach to improving the fairness of a model is to eliminate information about sensitive variables to prevent the model from using it (Kilbertus et al., 2017).
+This approach to fairness, which we’ll call "sensitive variable information erasure", has already been used in various forms in the literature
+One approach just removes explicit gender indicators (he, she, her, etc.) directly from the dataset (De-Arteaga et al., 2019).
+However, modern transformer models are able to learn implicit gender indicators (Devlin et al., 2018; Karita et al., 2019), which essentially limits the impact that such a naive approach can have on state-of-the-art models.
+In fact, a RoBERTa model trained without these gender indicators can still predict with upwards of 90% accuracy the gender of a person writing a short LinkedIn bio (see Appendix A).
+A second approach attempts to create word embeddings that do not exhibit gender bias (Bolukbasi et al., 2016; Garg et al., 2018; Zhao et al., 2018a, b; Caliskan et al., 2017; Bolukbasi et al., 2016; Zhao et al., 2019).
+Although interesting in principle, Gonen and Goldberg (2019) shows these methods may camouflage gender biases in word embeddings, but fail to eliminate them entirely.
+A third approach, more in line with our proposed method, consists in intervening on the model’s latent space.
+Liang et al.
+(2020) introduce a technique based on decomposing the latent space using a PCA and subtracting from the inputs the projection into a bias subspace that carries the most information about the target concept.
+Another vein focuses on modifying the latent space with considerations coming from the field of information theory (Xu et al., 2020).
+Indeed, the $\nu$-information – a generalization of the mutual information – has been used as a means to carefully remove from the model’s latent space only the information related to the target concept.
+In particular, popular approaches include the removal of information through linear adversarials on the last layer of the model (Ravfogel et al., 2020, 2022), through Assignment-Maximization adversarials (Shao et al., 2023), and through linear models that act on the whole model (Belrose et al., 2023).
+However, Ravfogel et al.
+(2023) has recently shown a limitation of these methods.
+The limitation comes from the practical use of the $\nu$-information ($\nu$-information is a metric that updates the mutual information, considering the expressivity of functions measuring information in the latent space).
+Ravfogel et al.
+(2023) explains that by protecting information using a linear model, more expressive models can still decode this protected information.
+Our method avoids this problem, because we are effectively destroying information – and not simply hiding it from linear models.
+Thus, no model can access it.
+In addition, our method, inspired by concept-based explainability techniques like TCAV (Kim et al., 2018), CRAFT (Fel et al., 2023), and COCKATIEL (Jourdan et al., 2023), provides the added benefit of enabling interpretability of the erased concepts.
+Unlike most concept erasure methods, which focus solely on removing sensitive information, our approach offers insight into how these concepts influence model predictions, enhancing transparency and interpretability without sacrificing performance.
+Figure 1: Overview of TaCo method.
+A decomposition of the final latent embedding matrix (I) yields concepts, whose importance (II) with respect to the sensitive variable and the label are evaluated with Sobol method.
+Finally, some concepts are removed, which “neutralizes” the sensitive variable information and (III) produces a fairer classifier.
 3 Preliminaries and definitions
-3.1 Methodology
-The idea behind the method we propose for creating gender-neutral embedding can be summed up in 3 components: (i) Identification of occupation-related concepts within the latent space of an NLP model; (ii) evaluation of the significance of each concept in predicting gender; and (iii) removal of specific concepts to construct a gender-neutral embedding, followed by retraining a classifier to predict occupation based on this modified embedding.
+The idea behind our method for concept erasure has 3 components: (i) identifies label-related concepts within the final latent space of an NLP model (such as the [CLS] embedding for encoder transformers like BERT, the EOS token embedding for encoder-decoder transformers like T5, or the first predicted token for decoder transformers like Llama).
+(ii) evaluates the importance of each concept for the prediction of the sensitive variable and the label to find the concepts to remove with the best trade-off between removing sensitive information and maintaining model performance; (iii) removes specific concepts to construct a new final latent embedding without sensitive variable information, after which we retrain a classifier to predict label based on this modified embedding.
 For a visual representation of the method, refer to Figure 1, which provides an explanatory diagram illustrating the different stages of the process.
-To realize the proposed method, it is crucial to secure certain guarantees and engage in rigorous theoretical modelling.
-This ensures the identification and application of suitable, practical tools.
-This involves establishing clear criteria for selecting practical tools and methodologies, enabling the achievement of dependable and consistent results.
-These are defined in this section.
+Our method relies on a mathematical foundation that guarantees removal and ensures the identification and application of suitable, practical tools.
+This foundation also establishes clear criteria for selecting practical tools and methodologies, enabling the achievement of dependable and consistent results.
+These are defined in the next section.
 3.1 Notation
-[…]
-With $n$ individuals in the training set, we define $\bm{X}=(\bm{x}_{1},...,\bm{x}_{n})\in\mathcal{X}^{n}$, $\bm{Y}=(y_{1},...,y_{n})\in\mathcal{Y}^{n}$, and $\bm{G}=(g_{1},...,g_{n})\in\{\text{Male, Female}\}^{n}$.
-The model $\bm{f}$ is conceptualized as a composition of two functions.
-First, $h$ encompasses all layers from the input $\bm{x}$ to a latent space, which is a transformed representation of the input data.
-Then, the classification layers $c$ classify these transformed data.
-The model can be summarized as:
+We work in context of supervised learning assume that a neural network model, denoted as $\bm{f}\colon\mathcal{X}\rightarrow\mathcal{Y}$, has been previously trained to perform a specific classification task.
+Here, $\bm{x}\in\mathcal{X}$ represents an embedded input text (in our example, a LinkedIn biography), and $y\in\mathcal{Y}$ denotes their corresponding label (in our example, the occupation).
+We introduce an additional set, $g\in\mathcal{G}$, representing the sensitive variables (in our example, the gender).
+This variable is crucial as it is the one we have to control to ensure fairness in our model.
+With $n$ individuals in the training set, we define $\bm{X}=(\bm{x}_{1},...,\bm{x}_{n})\in\mathcal{X}^{n}$, $\bm{Y}=(y_{1},...,y_{n})\in\mathcal{Y}^{n}$, and $\bm{G}=(g_{1},...,g_{n})\in\mathcal{G}^{n}$.
+We represent the model $\bm{f}$ as a composition of two functions.
 $\bm{f}(\bm{x})=c\circ h(\bm{x})\,,$ (1)
 where $h(\bm{x})\in\mathbb{R}^{d}$.
-In this space, we obtain an output embedding matrix, denoted as $\bm{A}=h(\bm{X})\in\mathbb{R}^{n\times d}$, which is also referred to as the CLS embedding in the context of an encoder transformer for classification task.
+$h$ encompasses all layers from the input $\bm{x}$ to a latent space, which is a transformed representation of the input data.
+The classification layers $c$ classify these transformed data.
+$h(\bm{X})$ is the final latent embedding matrix, $\bm{A}=h(\bm{X})\in\mathbb{R}^{n\times d}$, which is also referred to as the [CLS] embedding in the context of an encoder transformer, as the EOS token embedding for encoder-decoder transformers and the first predicted token for decoder transformers for a classification task.
 3.2 Causal underpinnings of the method
-A faithful reconstruction of $\bm{f}$ should support a causal relation that can be formulated as follows: $A$ and $B$ and had $A$ not been the case, $B$ would not have been the case either Lewis (1973); Pearl (2009); Jacovi and Goldberg (2020).
+A faithful reconstruction of $\bm{f}$ should support a causal relation: $A$ and $B$ and had $A$ not been the case, $B$ would not have been the case either (Lewis, 1973; Pearl, 2009; Jacovi and Goldberg, 2020).
 $A$ describes the causally necessary and “other things being equal” sufficient condition for $B$.
-Faithfulness implies that a particular type of input is causally necessary for the prediction at a given data point $\bm{x}$.
-Such a condition could take the form of one or more variables having certain values or set of variables.
-It can also take the form of a logical statement at a higher degree of abstraction.
-A counterfactual theory is simply a collection of such statements, and for every deep learning model, there is a counterfactual theory that describes it Jacovi et al.
-(2021); Asher et al.
-(2022); Yin and Neubig (2022).
+Faithfulness implies that a particular input type is causally necessary for the prediction at a given data point $\bm{x}$.
+Such a condition could involve specific variable values or a set of variables and can also be a higher-level logical statement.
 […]
-For instance, we can alternatively choose (i) $\mathscr{W}=\mathcal{X}^{n}$ the model’s input entries, (ii) $\mathscr{W}=$ the set of possible attention weights in $\bm{f}$ (assuming it’s a transformer model), (iii) $\mathscr{W}=$ the possible parameter settings for a specific intermediate layer of $\bm{f}$ –e.g., the attribution matrix as in Fel et al.
-(2023), (iv) $\mathscr{W}=$ the possible parameter settings for the final layer of an attention model as suggested in Wiegreffe and Pinter (2019), or (iv) sets of factors or dimensions of the final layer output matrix from the entire model Jourdan et al.
-(2023).
-Our explanatory model here follows the ideas of Jourdan et al.
-(2023), though it differs importantly in details.
-For one thing, we note that our general semantic approach allows us to have Boolean combinations even first order combinations of factors to isolate a causally necessary and sufficient explanans.
+For example, we can choose: (i) $\mathscr{W}=\mathcal{X}^{n}$ (model’s input entries), (ii) $\mathscr{W}$ as the set of possible attention weights in $\bm{f}$ (if it’s a transformer model), (iii) $\mathscr{W}$ as the possible parameter settings for a specific intermediate layer of $\bm{f}$ (e.g., attribution matrix as in (Fel et al., 2023)), (iv) $\mathscr{W}$ as the possible parameter settings for the final layer of an attention model as suggested in (Wiegreffe and Pinter, 2019), or (v) sets of factors or dimensions of the final layer embedding matrix from the entire model (Jourdan et al., 2023).
+Our explanatory model follows the ideas of (Jourdan et al., 2023) but differs in details.
+Notably, our semantic approach allows Boolean combinations, even first-order combinations of factors, to isolate a causally necessary and sufficient explanans.
 3.2.1 Counterfactual intervention for concept ranking
-In part 2 of our method, we calculate the causal effect on the sensitive variable for each dimension $r$ of $\bm{U}$ and $\bm{W}$ of the $\bm{A}=\bm{U}\bm{W}$ decomposition created in part 1 of the method (more details on this decomposition in practice in section 4.1).
-We conduct a counterfactual intervention that involves removing the dimension or “concept” $r$.
-By comparing the behaviour of the model with the concept removed and the behaviour of the original model, we can assess the causal effect of the concept on the sensitive variable.
+In part 2 of our method, we calculate the causal effect on the sensitive variable for each dimension $r$ of $\bm{U}$ and $\bm{W}$ of the $\bm{A}=\bm{U}\bm{W}$ decomposition created in part 1 of the method (for more details on this decomposition see section 4.1).
+We conduct a counterfactual intervention that involves removing the dimension or “concept” $k\in[\![1;r-1]\!]$.
+By comparing the behavior of the model with the concept removed and the behavior of the original model, we can assess the causal effect of the concept on the sensitive variable.
 As explained above, this assumes a counterfactual model where $\mathscr{W}$ corresponds to the possible variations of $\bm{A}=h(\bm{X})$ when reducing the dimensionality of its decomposition $\bm{A}=\bm{U}\bm{W}$.
-In addition, this approach requires that we are really able to remove the targeted dimension, which means that the dimensions must be independent of each other.
+To really able to remove the targeted dimension, the dimensions must be orthogonal of each other. Our mathematical framework guarantees this as far as possible.
 3.3 Information theory guarantees
-After performing our counterfactual intervention, we need to compute the "importance of a dimension" for the sensitive variable.
-Our goal is for our final embedding $\bm{A}$ to be as independent as possible from the sensitive variable $\bm{G}$.
-According to information theory, if we are unable to predict a variable $\bm{G}$ from the data $\bm{A}$, it implies that this variable $\bm{G}$ does not influence the prediction of $y$ by $\bm{A}$ Xu et al.
-(2020).
-This definition is widely used in causal fairness studies Kilbertus et al.
-(2017), making it a logical choice for our task.
-We train a classifier $c_{g}$ for the gender $\bm{G}$ prediction from the matrix $\bm{A}$ (see appendix C.2 for details on its architecture).
-If $c_{g}$ cannot achieve high accuracy for gender prediction, it implies that $\bm{A}$ does not contain gender-related information.
-The most reliable approach is thus to remove each of the dimensions found by the decomposition one by one and re-train a classifier to predict gender from this matrix.
-This can be computationally expensive if we have many dimensions.
-An alternative is to train only one classifier to predict gender on the original matrix $\bm{A}$, and then use a sensitivity analysis method on the dimensions to determine their importance for prediction.
-We will choose the latter for our method.
-However, we will still demonstrate the relevance of this approach by training a classifier each time we remove a dimension and show its effect in Section 5.2.
+As part of our counterfactual intervention, we calculate the "importance of a dimension" for the sensitive variable, so that our final integration $\bm{A}$ is as independent as possible from the sensitive variable $\bm{G}$.
+According to information theory, if we are unable to predict a variable $\bm{G}$ from the data $\bm{A}$, it implies that this variable $\bm{G}$ does not influence the prediction of $y$ by $\bm{A}$ (Xu et al., 2020).
+This definition is widely used in causal fairness studies (Kilbertus et al., 2017), making it a logical choice for our task.
+To evaluate the independence between $\mathbf{a}$ and $\mathbf{g}$, we train a classifier $c_{g}$ to predict the sensitive variable $\mathbf{g}$ from the representation $\mathbf{a}$ (see Appendix C.2 for architectural details).
+If $c_{g}$ cannot achieve high accuracy in predicting $\mathbf{g}$, it suggests that $\mathbf{a}$ lacks information about the sensitive attribute.
+While existing methods often employ linear classifiers to estimate this predictability, focusing solely on linear dependencies may overlook non-linear relationships, allowing residual sensitive information to remain accessible to non-linear models (Ravfogel et al., 2023).
+To address this limitation, we utilize a non-linear classifier—specifically, a multilayer perceptron (MLP)—to estimate the predictability of $\mathbf{g}$ from $\mathbf{a}$.
+This approach enables us to detect and mitigate both linear and non-linear dependencies, providing a more robust assessment of the residual sensitive information.
+Calculating the importance of each dimension by individually removing them and retraining the classifier can be computationally intensive, especially in high-dimensional spaces.
+To circumvent this, we train a single non-linear classifier on the original representation $\mathbf{a}$ and apply a sensitivity analysis method (Section 4.2) to determine each dimension’s importance for predicting $\mathbf{g}$.
+This strategy efficiently identifies the dimensions most influential to the sensitive attribute without the need for exhaustive retraining.
+For quantifying the degree of dependence between $\mathbf{a}$ and $\mathbf{g}$, we adopt the $\nu$-information metric.
+While this metric has been well-established in prior work using linear classifiers (Ravfogel et al., 2023; Belrose et al., 2023), we extend its application by incorporating non-linear classifiers in its computation.
+This extension allows us to capture non-linear dependencies that linear models might miss, offering a more comprehensive measure of mutual information.
+By using a non-linear classifier to compute $\nu$-information, we provide a stronger guarantee of fairness against adversaries employing complex, non-linear models.
+In Section 5.2, we show that our method significantly reduces the presence of sensitive information, as evidenced by the decreased predictability of $\mathbf{g}$ even when using sophisticated non-linear classifiers.
 4 Methodology
-Now that we’ve established the constraints on how to implement our method, we define the tools used for each part of the method: (i) For the decomposition that uncovers our concepts, we use the SVD decomposition.
-(ii) For the gender importance calculation, we train a classifier for this task and then use Sobol importance.
-(iii) To remove concepts, we simply remove the columns corresponding to the removed concept.
+Now that we’ve established constraints on how to implement our method, we define the tools used for each part of the method.
+(i) For the decomposition that uncovers our concepts, we use different decompositions: SVD, PCA, and ICA.
+(ii) For the sensitive variable importance calculation, we train a classifier for this task and then use Sobol importance.
+(iii) To remove concepts, we simply remove the columns of the matrix that correspond to the target concept.
 4.1 Generating concepts used for occupation classification
-Firstly, we look into latent space before the last layer of the model, where the decisions are linear.
-We have an output embedding matrix (also called CLS embedding) of the transformer model in this space.
-The aim here is to employ decomposition techniques to extract both a concept coefficients matrix and a concept base matrix from latent vectors.
-To illustrate this part, we discover concepts without supervision by factorizing the final embedding matrix through a Singular Value Decomposition.
-We choose this technique, but it is possible to do other decompositions, like PCA Wold et al.
-(1987), RCA Bingham and Mannila (2001) or ICA Comon (1994).
-Singular Value Decomposition (SVD)
-The embedding matrix $\bm{A}\in\mathbb{R}^{n\times d}$ can be decomposed using an SVD as:
-$\bm{A}=U_{0}\Sigma_{0}V_{0}^{\intercal}=U_{0}W_{0}\,,$
-where $U_{0}\in\mathbb{R}^{n\times d}$ contains the left singular vectors, $V_{0}\in\mathbb{R}^{d\times d}$ contains the right singular vectors, and $\Sigma_{0}\in\mathbb{R}^{d\times d}$ is a diagonal matrix containing the singular values.
-Note that $U_{0}$ and $V_{0}$ are orthonormal matrices, which means that $U_{0}^{\intercal}U_{0}=Id$ and $V_{0}^{\intercal}V_{0}=Id$.
-An SVD can be used to explain the main sources of variability of the decomposed matrix, here $\bm{A}$.
-The matrix $W_{0}^{-1}=V_{0}\Sigma_{0}^{-1}$ is indeed considered as a linear projector of $\bm{A}$, and $U_{0}$ is the projection of $\bm{A}$ on a basis defined by $V_{0}$.
-The amount of variability, or more generally, of information that has been captured in $\bm{A}$ by a column of $U_{0}$ is directly represented by the corresponding singular value in $\Sigma$.
-Using only the $r$ largest singular values of $\Sigma_{0}$, with $r<<d$, therefore allows capturing as much variability in $\bm{A}$ as possible with a reduced amount of dimensions.
-We then decompose $\bm{A}$ as:
+Firstly, we examine the latent space before the last layer of the model.
+We have an final latent embedding $\bm{A}\in\mathbb{R}^{n\times d}$ of the transformer model in this space.
+The aim here is to employ matrix decomposition techniques to extract both a concept coefficients matrix $\bm{U}$ and a concept base matrix $\bm{W}$ for the latent vectors.
+To illustrate this part, we discover concepts without supervision by factorizing the final embedding matrix using different decompositions: singular value decomposition (SVD), principal component analysis (PCA) (Wold et al., 1987) and independent component analysis (ICA) (Comon, 1994).
+For SVD, the matrix $\bm{A}$ is decomposed into $\bm{A}=U_{0}\Sigma_{0}V_{0}^{\intercal}$, where $U_{0}\in\mathbb{R}^{n\times n}$ and $V_{0}\in\mathbb{R}^{d\times d}$ are orthonormal matrices, and $\Sigma_{0}\in\mathbb{R}^{n\times d}$ is diagonal.
+This decomposition reveals the main variability in $\bm{A}$.
+By retaining the $r\ll d$ largest singular values in $\Sigma$, we approximate $\bm{A}$ as:
 $\bm{A}\approx\bm{U}\bm{W}\,,$ (2)
-where $\bm{W}=\Sigma V^{\intercal}$ is a compressed version of $W_{0}$, in the sense that $\Sigma\in\mathbb{R}^{r,r}$ is the diagonal matrix containing the $r$ largest singular values of $\Sigma$ and $V\in\mathbb{R}^{d,r}$ contains the $r$ corresponding right singular vectors.
-The matrix $\bm{U}\in\mathbb{R}^{n,r}$ consequently contains the $r$ right singular vectors which optimally represent the information contained in $\bm{A}$ according to the SVD decomposition.
-In this paper, we refer to these vectors as the extracted concepts of $\bm{A}$.
-Python implementation details are in Appendix D.
+where $\bm{U}\in\mathbb{R}^{n\times r}$ contains the leading $r$ columns of $U_{0}$, and $\bm{W}=\Sigma V^{\intercal}\in\mathbb{R}^{r\times d}$ combines the singular values and right singular vectors (see (Eckart and Young, 1936)).
+SVD is used to capture the most significant patterns in the data by reducing dimensionality while preserving maximum variance.
+In PCA and ICA, $\bm{A}$ is projected onto the directions of maximum variance, resulting in a similar approximation:
+$\bm{A}\approx\bm{U}\bm{W}\,,$
+where $\bm{U}\in\mathbb{R}^{n\times r}$ contains the top $r$ principal components (the eigenvectors of the covariance matrix of $\bm{A}$) for the PCA or contains $r$ independent components for the ICA, and $\bm{W}\in\mathbb{R}^{r\times d}$ contains the corresponding loadings for the PCA or is the mixing matrix for the ICA.
+We use PCA to reduce dimensionality by capturing as much variance as possible with the fewest components, making it useful for data compression and noise reduction.
+And ICA is used to identify underlying factors or sources in the data that are mutually independent, revealing hidden structures not accessible via PCA or SVD.
+By expressing $\bm{A}$ as a product of $\bm{U}$ and $\bm{W}$ across these methods, we obtain lower-dimensional representations that capture essential characteristics of the data while emphasizing different statistical properties: SVD focuses on capturing overall variability, PCA emphasizes variance and decorrelation, and ICA highlights statistical independence.
+Note that the ICA does not give orthogonal dimensions, as we requested in the decomposition prerequisites, but we decide to study it here anyway, in the hope that the statistical independence between its dimensions may bring interesting results.
+For more details on the implementation, we refer the reader to Appendix D.
 4.2 Estimating concept importance with Sobol Indices
@@
 We then evaluate the concepts’ importance by applying a feature importance technique on $\bm{G}$ using $c_{g}$, and on $\bm{Y}$ using $c$.
-More specifically, Sobol indices Sobol (1993) are used to estimate the concepts’ importance, as in Jourdan et al.
-(2023).
-A key novelty of the methodology developed in our paper is to simultaneously measure each concept importance for the outputs $\bm{G}$ and $\bm{Y}$, and not only $\bm{Y}$ as proposed in Jourdan et al.
-(2023).
-Figure 2: Co-importance plot for $r=20$ dimensions with respect to Occupation and Gender labels, computed with the Sobol method.
-The colour is based on the angle $a=90-\frac{2}{\pi}\arctan{(\frac{y}{x})}$.
-The angles with extreme values correspond to concepts with high importance for gender, but comparatively low importance for occupation.
-Their removal yields a favourable tradeoff in accuracy/fairness.
-To estimate the importance of a concept $\bm{U}_{i}$ for gender (resp.
-for the label), we measure the fluctuations of the model output $c_{g}(\bm{U}\bm{W})$ (resp.
-$c(\bm{U}\bm{W})$) in response to perturbations of concept coefficient $\bm{U}_{i}$
+More specifically, Sobol indices (Sobol, 1993) are used to estimate the concepts’ importance, as in (Jourdan et al., 2023).
+A key novelty of the methodology developed in our paper is to simultaneously measure each concept’s importance for the outputs $\bm{G}$ and $\bm{Y}$, and not only $\bm{Y}$ as proposed in (Jourdan et al., 2023).
+Figure 2: Co-importance plot for $r=20$ dimensions with respect to Occupation and Gender labels, computed with the Sobol method on RoBERTa model.
+The color is based on the angle $a=90-\frac{2}{\pi}\arctan{(\frac{y}{x})}$.
+The angles with extreme values correspond to concepts of high importance for gender but comparatively low importance for occupation.
+Their removal yields a favorable tradeoff in accuracy/fairness.
+To estimate the importance of a concept $\bm{U}_{i}$ for sensitive variable (resp.
+for the label), we measure the fluctuations of the model’s output $c_{g}(\bm{U}\bm{W})$ (resp.
+$c(\bm{U}\bm{W})$) in response to perturbations of concept coefficient $\bm{U}_{i}$.
 We then propagate this perturbed activation to the model output $\bm{G}=c_{g}(\tilde{\bm{A}})$ (resp.
 $\bm{Y}=c(\tilde{\bm{A}})$).
-We can capture the importance that a concept might have as a main effect – along with its interactions with other concepts – on the model’s output by calculating the expected variance that would remain if all the concepts except the i were to be fixed.
+We can capture the importance that a concept might have as a main effect – along with its interactions with other concepts – on the model’s output by calculating the expected variance that would remain if all the concepts except the $i$-th were to be fixed.
 This yields the general definition of the total Sobol indices.
 An important concept will have a large variance in the model’s output, while an unused concept will barely change it.
-We explain the total Sobol indices for each concept in details in Appendix E, along with our perturbation strategy.
+We explain the total Sobol indices for each concept in detail in Appendix E, along with our perturbation strategy.
 4.3 Neutralizing the embedding with importance-based concept removal
-In this final section, we propose to eliminate $k$ concepts that exhibit significant gender dependence within the final embedding matrix, with the objective of improving its gender-neutrality.
-This provides us with an improved embedding which can serve as a base to learn a fairer classifier to predict downstream tasks, and thus, a more gender-neutral model overall.
-This concept-removal procedure consists of choosing the $k$ concepts that maximize the ratio between the importance for gender prediction and for the task.
-This graphically corresponds to searching for the concepts whose angle with respect to the line of equal importance for gender and task prediction are the highest (see Fig.
+In this final section, we propose to eliminate $k$ concepts that exhibit significant dependence on sensitive variables from the final embedding matrix.
+This provides us with an improved embedding which can serve as a base to learn a fairer classifier to predict downstream tasks and, thus, a model that is globally more neutral with regard to sensitive variables.
+This concept-removal procedure consists of choosing the $k$ concepts that maximize the ratio between the importance for sensitive variable prediction and for the task.
+This graphically corresponds to searching for the concepts whose angle with respect to the line of equal importance for sensitive variable and task prediction are the highest (see Fig.
 2).
 This strategy allows us to work on a trade-off between the model’s performance and its fairness.
-In practice, we sort the concepts $\bm{U}_{1}$,$\bm{U}_{2}$,….$\bm{U}_{r}$ according to the second strategy in descending order and we delete the $k$ first concepts.
+In practice, we sort the concepts $\bm{U}_{1}$,$\bm{U}_{2}$,….$\bm{U}_{r}$ according to this strategy in descending order and delete the first $k$ concepts.
 This essentially leaves us with a concept base $\tilde{\bm{U}}=(\bm{U}_{k+1},...,\bm{U}_{r})$.
-This same transformation can be applied to $\bm{W}$’s lines, yielding $\tilde{\bm{W}}$.
+This same transformation can be applied to $\bm{W}$’s rows, yielding $\tilde{\bm{W}}$.
 From the matrix obtained from the matrix product $\tilde{\bm{U}}\tilde{\bm{W}}$, we retrain a classifier $\tilde{c}$ to predict $\bm{Y}$ (refer to Appendix C.2 for more details).
-Remark:
-The more concepts we remove, the less we will be able to predict the gender, so the fairer the model will be.
-However, this also entails a decrease on the amount of information that is encoded in the embedding and that can be used for downstream tasks.
-The parameter $k$ must therefore be chosen to comply with our specific application’s constraints.
-Figure 3: Accuracy drop when $k$ concepts are removed for an SVD with $r=20$ dimensions.
-In Figure 2(a), (Top) the occupation accuracy, (Bottom) the gender accuracy.
-Figure 2(b) shows occupation accuracy as a function of gender accuracy drop.
-The gender accuracy drop is reported relatively to the Baseline1, the classifier trained on Bios with all the concepts.
+It is interesting to note that, because we are causally removing information from the final latent space, it is impossible for any model to recover it from the truncated space – i.e.
+we are not limited by the family of models that’s intervening (Ravfogel et al., 2023).
+Figure 3: Gender prediction accuracy versus the occupation accuracy drop – using a two-layer MLP – after concept erasure methods on (top left) DistilBERT, (top right) RoBERTa, (bottom left) DeBERTa, and (bottom right) T5 representations.
+The occupation accuracy drop is reported relatively to the Baseline model, when the value is negative for a method, this means that occupation accuracy is better with the method than it was initially.
+For INLP and RLACE, points represent different numbers of masked dimensions.
+For TaCo methods, points represent different numbers of removed concepts.
+LEACE and the Bios-neutral model are shown as single points, as they do not require parameter adjustments.
+The horizontal dashed line represents the accuracy of the Optimal Bayes classifier, indicating the theoretical lower bound for gender prediction accuracy.
+Figures for linear classifier on Figure 7.
 5 Results
 5.1 Dataset and models
-To illustrate our method, we use the Bios dataset De-Arteaga et al.
-(2019), which contains about 440K biographies (textual data) with labels for the genders (M or F) of the authors, and their occupations (28 classes).
-From the Bios dataset, we also create a second dataset Bios-neutral by removing all the explicit gender indicators.
-Bios-neutral will be used as a baseline to which we can compare our method.
-We leave the details of how this dataset preprocessing was performed in Appendix B.
-For both datasets, we train a RoBERTa (Liu et al., 2019a) model to predict the occupation task. We leave the details on how these models were trained in Appendix C.
-5.2 Results analysis - Quantitative part
-Figure 4: Understanding concepts extracted from the model trained on Bios.
-Concept 1 prominently features explicit gender indicators like ’she’ and ’his’, crucial for gender identification but irrelevant for occupation.
-Concept 8 integrates information about gender-unbalanced professions, aiding in predicting both, an individual’s gender and occupation.
-Concept 19 focuses on information related to the professor occupation, a gender-balanced class and most represented in the dataset.
-Initially, we will examine the progression of our model’s performance for each dimension removed (one by one) utilizing our methodology.
-Following the precepts of information theory, we will assess the effectiveness of our bias mitigation strategy by measuring the model’s capacity to predict the gender and the task as we progressively remove concepts.
-We showcase in Fig.
-3 the outcome of removing the $k$ most important concepts for $k\in\llbracket 0,19\rrbracket$ (we have $r=20$ dimensions) in terms of different metrics: accuracy on gender prediction and on predicting occupation (downstream task).
-We compare our technique to two baselines: Baseline1 – RoBERTa model trained on Bios dataset (i.e.
-when $k=0$) –, and Baseline2 – RoBERTa model trained on Bios-neutral dataset.
-From Figure 3, simply by removing one concept, we achieve significant gains in the gender neutrality of our model, as it predicts gender with much lower accuracy (82.3% compared to the initial 96%) while maintaining its performance in occupation prediction (accuracy decreasing from 86.4% to 86.3%).
-Starting from the removal of the fourth dimension, the accuracy of gender prediction falls below the threshold of Baseline2, with 79.4% accuracy compared to 79.8% for Baseline2, while maintaining an accuracy of 86.3% for occupation (compared to 86% for Baseline2).
-Moreover, our method, in addition to being explainable, time-efficient, and cost-effective, exhibits greater effectiveness.
-After removing these first four concepts, we enter a regime where removing concepts continues to gain us fairness (as the accuracy of the gender task continues to fall), while the accuracy of the occupation remains unchanged.
-This regime is therefore the one with the best trade-off between accuracy and fairness (shown in green here).
-From concept 12 onwards, the accuracy of occupation begins to decline, as we look more and more at concepts with a high importance for occupation (according to Sobol), and we enter a new regime where we gain more and more in fairness, but where the accuracy of our task collapses.
-It may be worthwhile deleting a fairly large number of concepts to get into this zone, but only if we are willing to pay the price in terms of prediction quality.
-It is important to note that in this specific application, the optimal fairness-accuracy trade-off corresponds to accuracy on gender prediction of 62% as that’s the accuracy one gets when learning a model to predict the gender with the occupation labels as only information.
-5.3 Concepts removed analysis - Explainability part
-In this section, we interpret the dimensions of the Singular Value Decomposition (SVD) as valuable concepts for predicting our two tasks: gender and occupation.
-For this, we use the interpretability part of the COCKATIEL Jourdan et al.
-(2023) method on concepts/dimensions previously found with the SVD and ranked with the Sobol method by importance for gender and occupation.
-Concept interpretability
-COCKATIEL modifies the Occlusion method Zeiler and Fergus (2014), which operates by masking each word and subsequently observing the resultant impact on the model’s output.
-In this instance, to infer the significance of each word pertaining to a specific concept, words within a sentence are obscured, and the influence of the modified sentence (devoid of the words) on the concept is measured.
-This procedure can be executed at either the word or clause level, meaning it can obscure individual words or entire clauses, yielding explanations of varying granularity contingent upon the application.
-Concepts interpretation
-For the dimensions of the SVD to be meaningful for explanation, they must first be meaningful for the model’s prediction, whether it be the model predicting gender or the one predicting occupation.
-We rank the concepts w.r.t their removal order, and we focus on three important ones:
-Concept 1, with a gender importance of 32%, and an occupation importance of 5%; Concept 8, with a gender importance of 16%, and an occupation importance of 12%; Concept 19, with a gender importance of 12%, and an occupation importance of 33%.
-These three concepts are crucial for the prediction of at least one of our two tasks.
-Moreover, they are representative of our problem statement as Concept 1 is more important for gender, Concept 8 is as important to gender as occupation, and Concept 19 is more important for occupation.
-We illustrate several examples of explanations for these three concepts in Figure 4 to enable qualitative analysis.
-According to the importance per dimension calculated in part 2 of our method, concept 1 is almost exclusively predictive of gender.
-The most significant words for the concept of this dimension are explicit indicators of traditional gender.
-Concept 8 is capable of predicting both gender and occupation.
-Within its concept, we find examples of biographies predominantly related to unbalanced professions such as dietitian (92.8% women), chiropractor (26.8% women), software engineer (16.2% women), and model (83% women).
-It is logical that the concept allowing the prediction of these professions also enables the prediction of gender, as it is strongly correlated in these particular professions.
-Concept 19, which strongly predicts occupation, focuses on biographies predicting professors (constituting more than 29.7% of all individuals).
-This class is gender-balanced (with 45.1% women), providing little information on gender.
+To evaluate the efficacy of our method, we employ the Bios dataset (De-Arteaga et al., 2019), which comprises approximately 440,000 biographies annotated with the gender (male or female) of the individuals and their occupations across 28 classes.
+From this dataset, we derive a second dataset, Bios-neutral, by removing all explicit gender indicators.
+This Bios-neutral dataset serves as a baseline for comparison.
+The specifics of the dataset preprocessing are detailed in Appendix B.
+For both datasets, we train three encoder transformer models: RoBERTa (Liu et al., 2019a), DistilBERT (Sanh et al., 2019), and DeBERTa (He et al., 2020), and an encoder-decoder transformer model: T5 (Raffel et al., 2020), to perform the occupation prediction task.
+The training procedures and hyperparameter settings for these models are provided in Appendix C.1.
+5.2 Quantitative Evaluation of Concept Erasure
+We compare our method against several state-of-the-art concept erasure techniques:
+$\bullet$ Iterative Nullspace Projection (INLP) (Ravfogel et al., 2020)
+$\bullet$ Reduced Linear Adversarial Concept Erasure (RLACE) (Ravfogel et al., 2022)
+$\bullet$ Linear Adversarial Concept Erasure (LEACE) (Belrose et al., 2023).
+Additionally, we include:
+$\bullet$ Baseline Model: The original model without any modifications.
+$\bullet$ Bios-neutral Model: A model similar to the baseline but trained on the Bios-neutral dataset.
+$\bullet$ Optimal Bayes: corresponding to the optimal Bayes classifier $\mathbb{P}(G|Y)$ predicting gender from occupation.
+$\bullet$ TaCo Variants: Our proposed method using different decompositions—Singular Value Decomposition (SVD), Principal Component Analysis (PCA), and Independent Component Analysis (ICA).
+Figures 3 and 7 provide a comparative analysis of the debiasing methods, with the x-axis showing the drop in occupation prediction accuracy relative to the baseline model, and the y-axis representing the gender prediction accuracy.
+In these figures, the x-axis is truncated at a 0.1 accuracy drop, as losing more than 10% in task accuracy is not particularly informative.
+The full versions of these plots, without the x-axis truncation, are provided in Figures 8 and 9, on appendix.
+The varying points for each method correspond to different parameter settings:
+$\bullet$ INLP and RLACE: We manipulate the number of dimensions masked from the final latent representation (up to 768 dimensions for BERT-type representations and 512 dimensions for T5 representation).
+Specifically, for INLP, we masked [200, 300, 400, 500, 550, 600, 650, 700, 750] dimensions for BERT-type representations and [100, 200, 300, 350, 400, 450, 500] dimensions for T5 representation, and for RLACE, we masked [1, 5, 10, 100, 200, 500] dimensions for BERT-type representations and [1, 5, 10, 100, 200, 300] dimensions for T5 representation.
+RLACE typically requires fewer dimensions to be hidden to achieve effectiveness (Ravfogel et al., 2022).
+$\bullet$ TaCo Variants: For each decomposition (SVD, PCA, ICA), we vary the number of concepts (i.e., dimensions in the decomposition’s basis) removed, ranging from 1 to $r-1$, where $r=20$ is the number of components used.
+This results in 19 data points for each TaCo variant.
+$\bullet$ LEACE: Since LEACE does not require parameter tuning, it is represented by a single point.
+$\bullet$ Baseline and Bios-neutral Models: Each is depicted by a single point corresponding to their respective accuracies.
+$\bullet$ Optimal Bayes: Represented by a horizontal dashed line on the graph, as it has no associated occupation accuracy value.
+This line indicates the theoretical limit for gender prediction accuracy, below which further reduction is not meaningful.
+All results are averaged over five independent runs to ensure robustness.
+We assess the $\nu$-information in both linear and non-linear settings: linear $\nu$-information – measured by the accuracy of a logistic regression classifier predicting the sensitive attribute and the label –, and non-linear $\nu$-information – measured using a two-layer Multilayer Perceptron (MLP) classifier.
+Detailed results of these classifiers are provided in Appendix C.2.
+As illustrated in Figure 7 (in Appendix), LEACE and RLACE excel at obscuring sensitive information from linear classifiers, aligning with their design objectives.
+Notably, TaCo with PCA closely matches INLP’s performance in the linear context, despite INLP’s optimization for linear scenarios.
+In contrast, Figure 3 highlights TaCo’s superiority in non-linear settings.
+While LEACE and RLACE effectively conceal gender information from linear models, they fail to prevent non-linear models like MLPs from retrieving this information.
+TaCo, particularly with PCA and SVD decomposition, significantly reduces the gender information accessible to non-linear classifiers while maintaining high occupation prediction accuracy.
+This demonstrates a more favorable trade-off compared to simply removing explicit gender indicators, as done in the Bios-neutral model.
+Further analysis reveals that TaCo with PCA slightly outperforms TaCo with SVD in non-linear contexts, especially in scenarios where task accuracy remains largely unaffected.
+This suggests that PCA is the optimal decomposition for TaCo when aiming to balance task performance with sensitive information removal.
+Conversely, TaCo with ICA performs less effectively across both linear and non-linear settings.
+In summary, our proposed TaCo method, especially when combined with PCA decomposition, effectively mitigates the leakage of sensitive information in both linear and non-linear models.
+It achieves this with minimal impact on task performance, outperforming existing state-of-the-art techniques and offering a robust solution for bias mitigation in language models.
+5.3 Interpretability of Concepts Through Explainability Technique
+Beyond effectively erasing sensitive information from embeddings against non-linear adversaries, our method offers an additional advantage in terms of interpretability.
+Due to its structure, which parallels practices in explainable AI (XAI), TaCo facilitates the straightforward application of interpretability techniques to the identified concepts.
+Specifically, we applied the COCKATIEL method (Jourdan et al., 2023) to interpret the concepts derived from our matrix decompositions before their removal.
+This analysis, detailed in Appendix G, demonstrates that concepts obtained via SVD are semantically meaningful and provide insights into how the model processes information related to both the sensitive attribute and the primary task.
+However, the interpretability of these concepts varies according to the decomposition method used, so this must be taken into account when selecting the decomposition.
+This interpretability aspect underscores the transparency of our method and highlights its potential for enhancing model accountability, an advantage that may not be as readily achievable with other concept erasure techniques.
 6 Conclusion
-In this paper, we introduced a method for neutralizing the presence of bias in transformer model’s embeddings in an interpretable and cost-efficient way through concept-based explainability.
-In this study, we present a method applied to gender-related information, yet inherently adaptable to other sensitive variables.
-Indeed, by removing concepts that are mostly influential for gender prediction, we obtain embeddings that rely considerably less on gender information for generating predictions, and at the same time, remain highly accurate at the task.
-Furthermore, by doing so via explainable AI, we perform interventions that can be easily interpreted by humans, and thus gain trust in the bias mitigation process and insights into what produced the bias in the first place.
-Finally, thanks to its low computation cost, it can be easily applied to pre-trained, fine-tuned and even embeddings accessed through inference APIs, thus allowing our method to be applied to any model in a post-hoc manner.
+In this work, we introduced TaCo, a novel concept erasure method that removes sensitive information from latent space representations, ensuring fairness against both linear and non-linear adversaries.
+Our method involves concept discovery, ranking based on importance to both the sensitive attribute and target label, and removing concepts that affect the sensitive variable while preserving task performance.
+Experiments on the Bios dataset showed that TaCo outperforms state-of-the-art methods, effectively reducing non-linear classifiers’ ability to predict sensitive attributes without compromising model accuracy.
+Additionally, TaCo enhances interpretability by enabling insights into the factors influencing sensitive attribute predictions.
+By applying explainability techniques like COCKATIEL, we contribute to both model transparency and accountability.
+Our findings highlight the importance of addressing non-linear dependencies in mitigating biases, positioning TaCo as a significant step toward fairer and more transparent AI systems.
 Limitations
-Herein, we propose a novel approach that serves as a framework that can be applied to various concept extraction techniques and feature importance computation methods.
-This means that the SVD could be exchanged for other dimensionality reduction techniques (and thus, change the properties of the concept base and coefficients), or Shapley values could be preferred over Sobol indices for the relaxation on the independence assumption, for istance.
-In addition, our application was gender bias mitigation, but it could be applied to any other type of bias as long as the sensitive variable is available or can be estimated.
+While our proposed method demonstrates effectiveness in mitigating the influence of sensitive attributes, it does not guarantee a clear separation between concepts containing information about the output variable and those related to the sensitive attribute.
+In some cases, the concepts identified through decomposition may be inherently entangled, simultaneously encoding information pertinent to both the task and the sensitive variable.
+This entanglement poses a challenge in precisely isolating and removing only the unwanted information without adversely affecting the model’s performance on the primary task.
+Moreover, our approach to interpreting concepts post-removal has room for improvement.
+Interpreting concepts in natural language processing remains a complex endeavor due to the intricate and high-dimensional nature of textual data.
+Developing more sophisticated methods for concept interpretation is essential and warrants dedicated future research.
+Another limitation pertains to the scope of our application.
+While we focused on mitigating gender bias using the Bios dataset, which considers gender as a binary attribute, we acknowledge that gender is inherently non-binary.
+Consequently, models trained on such datasets may not fully capture the spectrum of gender identities present in real-world scenarios.
+Although our method is designed to operate effectively with variables encompassing multiple classes, including non-binary gender categorizations, the lack of datasets reflecting this diversity limits the evaluation of our approach in more inclusive settings.
+This underscores the need for the NLP community to develop and adopt datasets that incorporate non-binary and diverse representations of sensitive attributes, thereby promoting fairness and inclusivity in AI applications.
+Ethics Statement
+With the advent of LLMs and their expanding applications, it is crucial to ensure that biases are adequately addressed before deploying these models in high-stakes applications.
+In this paper, we introduce a method that not only mitigates biases but does so in a manner that promotes human understanding of the specific biases removed from the model.
+Transparency in identifying bias sources and understanding the trade-off involved are vital aspects of our approach.
+An important consideration is the computational efficiency of our method, which significantly reduces the resource burden.
+Our approach only requires training a classification head while keeping the backbone of the model unchanged.
+This computational affordability makes our method particularly valuable for companies with limited budgets, enabling them to effectively address bias concerns without incurring excessive costs.
+We believe that it is our ethical responsibility to prioritize fairness and accountability in the development and deployment of AI systems.
+By providing a debiasing technique that offers interpretability, computational efficiency, cost-effectiveness and ease of use, we aim to support the responsible use of LLMs, ensuring they are better aligned with ethical standards and advancing the progress of unbiased AI technologies.
 Acknowledgements
@@
 Appendix C Models details and convergence curves
-C.1 RoBERTa training
+C.1 Transformer training
 RoBERTa base training
-We use a RoBERTa model (Liu et al., 2019a), which is based on the transformers architecture and is pretrained with the Masked language modelling (MLM) objective.
-We specifically used a RoBERTa base model pretrained by Hugging Face.
+We use a RoBERTa model (Liu et al., 2019a), which is based on the transformers architecture and is pre-trained with the Masked language modeling (MLM) objective.
+We specifically used a RoBERTa base model pre-trained by HuggingFace.
 All information related to how it was trained can be found in Liu et al.
 (2019b).
-It can be remarked, that a very large training dataset was used to pretrain the model, as it was composed of five datasets: BookCorpus Zhu et al.
-(2015), a dataset containing 11,038 unpublished books; English Wikipedia (excluding lists, tables and headers); CC-News Mackenzie et al.
-(2020) which contains 63 millions English news articles crawled between September 2016 and February 2019; OpenWebText Radford et al.
-(2019) an open-source recreation of the WebText dataset used to train GPT-2; Stories Trinh and Le (2018) a dataset containing a subset of CommonCrawl data filtered to match the story-like style of Winograd schemas.
+It can be remarked, that a very large training dataset was used to pre-train the model, as it was composed of five datasets: BookCorpus (Zhu et al., 2015), a dataset containing 11,038 unpublished books; English Wikipedia (excluding lists, tables and headers); CC-News (Mackenzie et al., 2020) which contains 63 millions English news articles crawled between September 2016 and February 2019; OpenWebText (Radford et al., 2019) an open-source recreation of the WebText dataset used to train GPT-2; Stories (Trinh and Le, 2018) a dataset containing a subset of CommonCrawl data filtered to match the story-like style of Winograd schemas.
 Pre-training was performed on these data by randomly masking 15% of the words in each of the input sentences and then trying to predict the masked words.
-Figure 8: Convergence curves during training for occupation classification task, 15 epochs.
-Baseline1: RoBERTa model trained on Bios; Baseline2: RoBERTa model trained on Bios without explicit gender indicators.
-(On the left) accuracy curves; (on the right) loss curves.
+DistilBERT base training
+DistilBERT (Sanh et al., 2019) is a transformer architecture derivative from but smaller and faster than the original BERT (Devlin et al., 2018).
+This model is commonly used to do text classification.
+DistilBERT is trained on BookCorpus Zhu et al.
+(2015) (like BERT), a dataset consisting of 11,038 unpublished books and English Wikipedia (excluding lists, tables and headers), using the BERT base model as a teacher.
+DeBERTa base training
+DeBERTa (He et al., 2020) enhances the BERT and RoBERTa models by incorporating disentangled attention and an improved mask decoder.
+These enhancements enable DeBERTa to surpass RoBERTa in most Natural Language Understanding (NLU) tasks, utilizing 80GB of training data.
+In the DeBERTa V3 iteration (used here), DeBERTa is optimized efficiently through ELECTRA-Style pre-training and Gradient Disentangled Embedding Sharing.
+This version, compared to its predecessor, shows significant performance gains in downstream tasks.
+The base model of DeBERTa V3 features 12 layers with a hidden size of 768.
+It possesses 86M core parameters, and its vocabulary includes 128K tokens, adding 98M parameters to the Embedding layer.
+This model was trained on 160GB of data, similar to DeBERTa V2.
+T5-small base training
+The T5 model (Raffel et al., 2020), short for "Text-To-Text Transfer Transformer," follows a unified framework where all NLP tasks are cast as a text-to-text format, with both input and output represented as strings.
+This model, pre-trained by Hugging Face, is based on the transformer architecture and uses a sequence-to-sequence structure.
+T5-small, a more lightweight version of the full model, contains 60 million parameters and is pre-trained on the C4 dataset, a massive corpus extracted from publicly available web pages.
+The training objective involved a masked span prediction task, where consecutive tokens in the input are masked, and the model is trained to reconstruct the missing spans.
+This architecture and training objective allow T5 to perform a wide range of downstream tasks, including text classification, summarization, and translation, by leveraging its ability to convert any task into a text generation problem.
 Occupation prediction task
-After pre-training RoBERTa parameters on this huge dataset, we then trained it on the 400.000 biographies of the Bios dataset.
-The training was performed with PyTorch on 2 GPUs (Nvidia Quadro RTX6000 24GB RAM) for 10 epochs (we train 15 epochs but we stop at 10 epochs) with a batch size of 8 observations and a sequence length of 512 words.
+After pre-training RoBERTa/DistilBERT/DeBERTa/T5 parameters on this huge dataset, we then trained it on the 400.000 biographies of the Bios dataset.
+The training was performed with PyTorch on 2 GPUs (Nvidia Quadro RTX6000 24GB RAM) for 10/3/3/2 epochs (we train 15 epochs, but we perform early stopping, giving the best results at 10/3/3/2 epochs) with a batch size of 8 observations and a sequence length of 512 words.
 The optimizer was Adam with a learning rate of 1e-6, $\beta_{1}=0.9$, $\beta_{2}=0.98$, and $\epsilon=1e6$.
-We split dataset into 70% for training, 10% for validation, and 20% for testing.
-The convergence curves of the model are in Figure 8.
+We split the dataset into 70% for training, 10% for validation, and 20% for testing.
 C.2 Classifier heads training
+C.2.1 Non-linear classifiers
 The importance estimation in part 2 of our method requires a classifier to be trained on top of the features.
-An ideal classifier will leverage the maximum amount possible of information about gender or occupation that can be realistically extracted from latent features, in the spirit of the maximum information that can be extracted under computational constraints Xu et al.
-(2020).
+An ideal classifier will leverage the maximum amount possible of information about gender or occupation that can be realistically extracted from latent features, in the spirit of the maximum information that can be extracted under computational constraints (Xu et al., 2020).
 Hence, it is crucial to optimize an expressive model to the highest possible performance.
@@
 Note that the paper reports the accuracy on the test set.
-The gender classifier $c_{g}$ and the occupation classifer $\tilde{c}$ are both trained with this same protocol.
-Appendix D Implementation details for SVD
+The gender classifier $c_{g}$ and the occupation classifier $\tilde{c}$ are both trained with this same protocol.
+C.2.2 Linear classifiers
+To compare with state-of-the-art models, we also train simple logistic regressions for the gender classifier $c_{g}$ and the occupation classifier $\tilde{c}$.
+We use linear_model.LogisticRegression from the scikit-learn package.
+We don’t optimize specific parameters.
+Appendix D Implementation details for decompositions
 We used the SVD implementation in sparse.linalg.svds provided in the Scipy package.
@@
 Note that this implementation turned out to obtain the most robust decompositions of $\bm{A}$ after an empirical comparison between torch.svd, torch.linalg.svd, scipy.linalg.svd, and sklearn.decomposition.TruncatedSVD on our data.
+We used PCA and ICA implementations in decomposition.PCA and decomposition.FastICA respectivly provided in the scikit-learn package.
 Appendix E Sobol Indices - details
@@
 (2023), with the exception of changes to the $\phi$ function created here to manage concepts found in all the classes at the same time (unlike COCKATIEL, which concentrated on finding concepts in a per-class basis).
+Figure 7: Gender prediction accuracy versus the occupation accuracy drop – using a Logistic Regression – after concept erasure methods on (top left) DistilBERT, (top right) RoBERTa, (bottom left) DeBERTa, and (bottom right) T5 representations.
+The occupation accuracy drop is reported relatively to the Baseline model, when the value is negative for a method, this means that occupation accuracy is better with the method than it was initially.
+For INLP and RLACE, points represent different numbers of masked dimensions.
+For TaCo methods, points represent different numbers of removed concepts.
+LEACE and the Bios-neutral model are shown as single points, as they do not require parameter adjustments.
+The horizontal dashed line represents the accuracy of the Optimal Bayes classifier, indicating the theoretical lower bound for gender prediction accuracy.
+Figure 8: Gender prediction accuracy versus the occupation accuracy drop – using a two-layer MLP – after concept erasure methods on (top left) DistilBERT, (top right) RoBERTa, (bottom left) DeBERTa, and (bottom right) T5 representations.
+The occupation accuracy drop is reported relatively to the Baseline model, when the value is negative for a method, this means that occupation accuracy is better with the method than it was initially.
+For INLP and RLACE, points represent different numbers of masked dimensions.
+For TaCo methods, points represent different numbers of removed concepts.
+LEACE and the Bios-neutral model are shown as single points, as they do not require parameter adjustments.
+The horizontal dashed line represents the accuracy of the Optimal Bayes classifier, indicating the theoretical lower bound for gender prediction accuracy.
+Figure 9: Gender prediction accuracy versus the occupation accuracy drop – using a Logistic Regression – after concept erasure methods on (top left) DistilBERT, (top right) RoBERTa, (bottom left) DeBERTa, and (bottom right) T5 representations.
+The occupation accuracy drop is reported relatively to the Baseline model, when the value is negative for a method, this means that occupation accuracy is better with the method than it was initially.
+For INLP and RLACE, points represent different numbers of masked dimensions.
+For TaCo methods, points represent different numbers of removed concepts.
+LEACE and the Bios-neutral model are shown as single points, as they do not require parameter adjustments.
+The horizontal dashed line represents the accuracy of the Optimal Bayes classifier, indicating the theoretical lower bound for gender prediction accuracy.
+Appendix F Additional Figures
+Figure 7 presents the results for the linear case, complementing the non-linear results shown in the main paper, thus providing a comprehensive comparison of performance based on the nature of the classifiers used.
+Figures 8 and 9 show the full comparative analysis of the debiasing methods, with the x-axis depicting the drop in occupation prediction accuracy relative to the baseline model and the y-axis representing the gender prediction accuracy.
+Figure 10: Clear interpretability of SVD-Derived Concepts in RoBERTa Using COCKATIEL, showing distinct and meaningful concepts for gender and occupation.
+Concept 1 prominently features explicit gender indicators like ’she’ and ’his’, crucial for gender identification but irrelevant for occupation.
+Concept 8 integrates information about gender-imbalanced professions, aiding in predicting both, an individual’s gender and occupation.
+Concept 19 focuses on information related to the professor occupation, a gender-balanced class and most represented in the dataset.
+Other explanations for DistilBERT model on Figure 11.
+Figure 11: Clear interpretability of SVD-Derived Concepts in DistilBERT Using COCKATIEL, with identifiable and interpretable concepts for gender and occupation..
+Concept 1 prominently features explicit gender indicators like ’she’ and ’his’, crucial for gender identification but irrelevant for occupation, and some information on specific professions that are particularly gender imbalanced.
+Concept 1O integrates information about gender-imbalanced professions, aiding in predicting both, an individual’s gender and occupation, and also some information on occupation in general.
+Concept 18 focuses on information related to the occupation in general.
+Appendix G Interpretability of Concepts Through Explainability Techniques
+While the primary focus of our work is on concept erasure for non-linear problems, an additional advantage of our method is its potential to enhance interpretability and transparency in models.
+This benefit arises from the ability to interpret the concepts identified through various matrix decompositions before they are removed.
+Our method’s construction inherently facilitates this interpretability due to its similarity with explainable AI (XAI) techniques.
+Such an approach may not be as straightforward with other concept erasure methods.
+In this section, we explore this aspect by analyzing the interpretability of concepts derived from Singular Value Decomposition (SVD) and Principal Component Analysis (PCA) across different models representation.
+Our analysis is facilitated by applying the COCKATIEL method (Jourdan et al., 2023), which provides explanations for the identified concepts.
+G.1 Common Interpretability Procedure with COCKATIEL
+COCKATIEL modifies the Occlusion method (Zeiler and Fergus, 2014) by masking each word in a sentence and observing the resultant impact on the model’s output.
+This approach allows us to infer the significance of each word with respect to a specific concept.
+By executing this procedure at either the word or clause level, we obtain explanations of varying granularity, enhancing our understanding of how certain concepts influence model predictions.
+Importantly, this interpretability procedure is uniformly applied across all decompositions and models, enabling a consistent analysis of concepts derived from different methods and architectures.
+Our method’s structure, which involves decomposing representations into interpretable concepts and ranking them based on their importance to the sensitive attribute and the target label, aligns closely with practices in XAI.
+This alignment allows us to seamlessly apply interpretability techniques like COCKATIEL to our concepts.
+Such ease of explanation may not be readily achievable with other concept erasure methods that do not decompose the representation space in a manner conducive to concept-level interpretation.
+G.2 Interpretability of SVD-Derived Concepts
+Using SVD, we decomposed the latent representations and ranked the resulting concepts based on their importance for predicting the sensitive variable (gender) and the target label (occupation) using the Sobol method.
+We focused on three key concepts for each model, RoBERTa and DistilBERT, that were representative of different importance levels.
+RoBERTa Model Analysis
+  • Concept 1: High importance for gender (32%), low for occupation (5%).
+  • Concept 8: Moderate importance for both gender (16%) and occupation (12%).
+  • Concept 19: Low importance for gender (12%), high for occupation (33%).
+As illustrated in Figure 10, applying COCKATIEL to these concepts in the RoBERTa model revealed that:
+  • Concept 1 contained explicit gender indicators such as pronouns "he" and "she," making it highly predictive of gender.
+  • Concept 8 included information from biographies of gender-imbalanced professions (e.g., dietitian, software engineer), thus linking gender and occupation.
+  • Concept 19 focused on the profession of professors—a gender-balanced and highly represented occupation—contributing primarily to occupation prediction without significant gender bias.
+DistilBERT Model Analysis
+  • Concept 1: High importance for gender (36.51%), low for occupation (6.63%).
+  • Concept 10: Moderate importance for both gender (12.42%) and occupation (13.57%).
+  • Concept 18: Lower importance for gender (16.63%), high for occupation (37.45%).
+Figure 11 highlights similar findings with the DistilBERT model, although the distinctions between concepts were less pronounced. Nevertheless, the concepts still contained meaningful information:
+  • Concept 1 had explicit gender indicators and terms related to gender-imbalanced occupations.
+  • Concept 10 was informative for both gender and occupation, containing terms from both gender-balanced and imbalanced professions.
+  • Concept 18 primarily focused on occupation-related terms from frequently occurring professions in the dataset.
+These findings demonstrate that the concepts derived from SVD are interpretable and provide insights into how the model processes information related to gender and occupation.
+The ability to extract such explanations is facilitated by our method’s structure, which is designed to isolate and rank concepts based on their relevance to the sensitive attribute and the target task.
+This structural design, mirroring XAI techniques, enables straightforward application of interpretability methods like COCKATIEL.
+Figure 12: Poor interpretability of PCA-Derived Concepts in RoBERTa Using COCKATIEL, revealing less coherent and less meaningful concepts compared to SVD.
+G.3 Challenges with PCA-Derived Concepts
+The explanations generated by COCKATIEL did not yield coherent or meaningful insights.
+The concepts lacked clear semantic associations with either gender or occupation, making it difficult to understand their influence on model predictions (see Figure 12 for an example on the RoBERTa model).
+Despite PCA showing better performance in concept erasure effectiveness, the reduced interpretability of its components presents a challenge.
+This suggests that while PCA effectively removes sensitive information, the resulting components may not align with human-interpretable concepts.
+Our method’s structure, which emphasizes concept identification and ranking, seems to facilitate interpretability more effectively with decompositions like SVD.
+It’s important to note that this observation is based on the application of COCKATIEL.
+Other explainability methods might offer different perspectives and could potentially uncover meaningful interpretations of PCA-derived concepts.
+However, the ease with which we can apply COCKATIEL to our method, due to its construction that mirrors XAI practices, may not extend to other concept erasure methods that lack this alignment.
+G.4 Implications for Explainability
+The ability to interpret concepts is an additional advantage of our method, providing benefits for model transparency and accountability.
+The ease with which we can obtain explanations for the concepts is linked to the construction of our method, which is similar to XAI techniques.
+By decomposing the latent space into ranked concepts, we create a framework that naturally lends itself to interpretability.
+This structural alignment allows for straightforward application of interpretability tools like COCKATIEL, which might not be as easily applied to other concept erasure methods that do not decompose representations in a way that isolates individual concepts.
+As a result, our method not only mitigates the influence of sensitive attributes but also enhances the explainability of the model, offering insights into which concepts are being removed and why.
+Our findings suggest that the choice of decomposition and the method’s design can impact both the effectiveness of concept erasure and the interpretability of the resulting concepts.
+The SVD decomposition, combined with our method’s structure, facilitates the extraction of meaningful explanations, enhancing transparency.
+Given that interpretability is an added advantage rather than the central focus of our work, these findings open avenues for future research.
+Exploring alternative explainability techniques may enhance the interpretability of PCA-derived concepts or other decompositions.
+Additionally, investigating other decomposition methods or hybrid approaches could balance the trade-off between concept erasure effectiveness and interpretability.
+Our method’s structure, which aligns closely with explainability frameworks, allows for such integrations.
+By facilitating the examination of internal model representations, we contribute to the broader goals of developing AI systems that are not only fairer but also more transparent and interpretable.
+This added interpretability dimension is a unique strength of our approach and may not be as easily replicated in other concept erasure methods that lack a similar structural design.